### v1

# Cracking Vector Search Indexes

Abstract.
Retrieval Augmented Generation (RAG) uses vector databases to expand the expertise of an LLM model without having to retrain it. This idea can be applied over data lakes, leading to the notion of embeddings data lakes, i.e., a pool of vector databases ready to be used by RAGs. The key component in these systems is the indexes enabling Approximated Nearest Neighbor Search (ANNS). However, in data lakes, one cannot realistically expect to build indexes for every possible dataset. In this paper, we propose an adaptive, partition-based index, CrackIVF, that performs much better than up-front index building. CrackIVF starts answering queries by near brute force search and only expands as it sees enough queries. It does so by progressively adapting the index to the query workload. That way, queries can be answered right away without having to build a full index first. After seeing enough queries, CrackIVF will produce an index comparable to the best of those built using conventional techniques. As the experimental evaluation shows, CrackIVF can often answer more than 1 million queries before other approaches have even built the index and can start answering queries immediately, achieving 10-1000x faster initialization times. This makes it ideal when working with cold data or infrequently used data or as a way to bootstrap access to unseen datasets.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
1. Introduction
External data sources can complement Large language models (LLMs) (Brown et al., 2020) through the use of retrieval-augmented generation (RAG) (Lewis et al., 2020). In RAG, external data is represented as vector embeddings (i.e., learned representations of data that result in a multidimensional vector). The embeddings are indexed by a variety of means and searched through approximate nearest neighbor (ANN) search so that LLMs can answer queries on information they were not explicitly trained on. The effectiveness of RAG is tied to that of the ANN index structures, making them a critical component.
LLMs have been integrated with search engines (OpenAI, 2024) (e.g., Google, Bing, and Baidu). However, the vast repository of private, unstructured business data remains largely untapped. It has been estimated that unstructured data constitutes up to 80-90% of the total data volume (Merrill Lynch, 1998; MIT Sloan School of Management, 2021) and is often stored in open data formats on data lakes (Miller, 2018). However, as much as 70% of this data (Data Science Association, 2016) remains unused, classified as “dark” data (Heidorn, 2008; Zhang et al., 2016).
This data can be made discoverable (Chapman et al., 2020), but even then, the actual data often remains unindexed. For instance, as of March 2020, Google Research’s Dataset Search (Brickley et al., 2019) had indexed around 28 million structured and unstructured datasets (Benjelloun et al., 2020) based on their metadata. However, embedding-based vector search and RAG techniques make the promise to go a step further by enabling direct retrieval and question answering on the underlying data itself, inside of each dataset. This requires embedding the data, and the creation of approximate nearest neighbor (ANN) indexes for every single dataset, exposing the data to RAG systems. In this paper, we name such an approach embedding data lakes (EDL), where unstructured data is stored alongside its vector representations.
Efforts in this direction are already ongoing in both industry and academia. Databricks has deployed vector search capabilities (Databricks, 2024) over their Lakehouse (Armbrust et al., 2021), and new custom storage formats like Lance (LanceDB Developers, 2024) are being used to store embeddings on data lakes. Researchers are also directly exploring RAG-based techniques to query multi-modal data lakes (Tang et al., [n.d.]; Chen et al., 2023), as well as answering queries directly over unstructured data (Liu et al., 2025; Anderson et al., 2024). There is also work to directly define LLM-based data discovery systems over data lakes (An et al., 2025).
The issue we address in this paper is crucial to achieve the promise of embedding lakes (EDL): index selection at scale. EDL, in principle, would require building millions of ANN indexes across diverse datasets, modalities, workloads, and embeddings, making static index selection and pre-computation infeasible. In data lakes, some datasets may receive millions of queries, while others are rarely accessed. When selecting an index for embedding lakes, there is a trade-off. On the one hand, building an index requires significant upfront costs that might not pay off if the data is rarely used. On the other hand, skipping the indexing process and relying on a brute-force search is not scalable. Hence, data must be indexed using Approximate Nearest Neighbor (ANN) indexes but this raises two questions: when to index the data and how to index it.
Figure 1. Total time to answer the queries submitted for different indexing strategies vs the number of queries submitted
An optimal ANN index depends on workload patterns and future usage, which are often unknown. As shown in Figure 1, the index that minimizes total cumulative time spent on building and searching varies with the query volume. Larger indexes take longer to build and require many queries to justify their upfront cost (e.g., the large IVFlat configurations (5000, 16000 clusters) in the figure). Brute-force search (as shown in the figure), and smaller indexes (IVFlat 1000 clusters), minimize time-to-query for fast discovery but struggle at scale due to their higher query response times.
To tackle index selection at scale, we adopt a well-established strategy: carefully avoiding the need to build the full index upfront. In this, we draw inspiration from deferred data structuring (Karp et al., 1988) and database cracking (Idreos et al., 2007; Halim et al., 2012; Schuhknecht et al., 2016), which have been successfully applied to single-dimensional database indexing. Similar approaches have been extended to multidimensional data (Holanda et al., 2018; Zardbani et al., 2023; Pavlovic et al., 2018; Nerone et al., 2021). However, prior work does not scale to the embedding dimensionalities nor supports the k-Nearest Neighbour (k-NN) queries typical in RAG systems. Recently, the AV-Tree was the first adaptive index for high-dimensional k-Nearest Neighbor (k-NN) search (Lampropoulos et al., 2023). AV-Tree targets short-lived data, with workloads of up to one thousand queries, and is a tree-based Exact Nearest Neighbor (ENN) adaptive index. This makes it impractical for RAG-scale workloads, which rely on Approximate Nearest Neighbor (ANN) index structures to trade off a bit of accuracy for large performance gains. Nonetheless, the idea of dynamically building an index tailored to the workload remains compelling. In this paper, we demonstrate how to achieve this efficiently in the context of RAG and employing the same data structures (Sivic and Zisserman, 2003) and systems (Douze et al., 2024) used for ANN search workloads (Aumüller et al., 2020; Simhadri et al., 2022) rather than analytics over cold data.
We introduce CrackIVF, an incrementally built, partition-based ANN index that dynamically improves itself to adapt to increasing workload demands, minimizing both time-to-query and cumulative cost associated with suboptimal static index selection. CrackIVF evolves as a side effect of query execution. Each query is a candidate to add a new partition, or crack, while its search region is a candidate for local refinement, allowing the index to grow and improve as more queries arrive. Efficiency is maintained through two controls: one deciding where to apply cracking and refinement in the search space, and another controlling when, to balance indexing and search. In the paper we focus on a static setting (i.e., no data or query distribution shifts) and evaluate CrackIVF on multiple standard open source datasets for ANN search (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016).
For partition-based indices, index selection means deciding upfront how many partitions the index will use, which has an impact on the ANN search performance. This decision has to be made without knowing the future query workload. CrackIVF waits to build the index until it sees enough queries and decides on the clusters based on the query distribution over the search space. As more queries arrive, the clusters are redefined, and their number is increased to adapt to the increasing query workload. Asymptotically, CrackIVF builds an index that converges to similar to those pre-built, but it has been able to answer queries along the way.
Across benchmarks, CrackIVF consistently outperforms pre-built partition-based indexes. It converges to near-optimal query response time, yet achieves several orders of magnitude lower startup time than other indexes. It efficiently scales to large workloads and, in some cases, it can process 1 million queries before the baseline indexes have finished building. It is the only index that consistently remains near the Pareto frontier of minimum cumulative time across the number of queries posed to the system.
2. Background
In this section, we briefly cover the basics of approximated nearest neighbor search and IVF indexes.
Nearest Neighbor Search In the k-nearest neighbor (k-NN) search problem, we are given a set of points $P\in\mathbb{R}^{|P|\times d}$ and a query $q\in\mathbb{R}^{d}$. The aim is to find the $k$ points in $P$ that are closest to $q$ under some notion of distance or similarity (e.g., minimize the Euclidean distance (L2) or maximize the inner product (IP) to measure closeness). A straightforward Exact Nearest Neighbor (ENN) approach is a linear scan of every point, a.k.a brute-force search, incurring in $\mathcal{O}(|P|\cdot d)$ complexity. There exist more sophisticated exact indexing structures (e.g., KD-trees (Bentley, 1975), R-trees (Guttman, 1984), or VP-trees (Yianilos, 1993)). However, they often degrade to near-linear-scan performance in higher dimensions, due to the curse of dimensionality (Weber et al., 1998), making them impractical when $|P|$ or $d$ are large. With embeddings of hundreds to thousands of dimensions  (OpenAI, 2023) and real-world datasets at the scale of billions of data points (Simhadri et al., 2022), RAG systems turn to Approximate Nearest Neighbor (ANN) search techniques. They allow a small retrieval error $\varepsilon$ in exchange for better runtime or memory usage. This approximation quality can also be evaluated by the fraction of exact top-$k$ vectors recovered.
(a) Response Time with increasing Partitions and training K-Means.
(b) Queries exhibit a skewed access pattern over the search space that is often far from uniform.
(c) Finer partitioning of the Vector Space based on queries received.
Figure 2. Observations behind the design: (a) separating number of partitions and refinement of centroids; (b) access to the index is generally skewed towards certain regions; (c) regions queried often can be clustered and refined at a much lower granularity
Partition based IVF methods One of the most common ANN search structures is partition-based IVF indexes (Sivic and Zisserman, 2003; Chen et al., 2021; Sun et al., 2024). Instead of scanning all points in $P\subset\mathbb{R}^{d}$, they divide the vector space into $n_{\text{list}}$ disjoint groups, or partitions. At query time, the search process first identifies the $n_{\text{probe}}\leq n_{\text{list}}$ nearest partitions to $q$, whose representative vectors are closest (or most similar) to the query. Once these partitions are selected, the points within them are exhaustively scanned. Since the remaining partitions are skipped, some nearest neighbors may be missed, which is why the result is an approximate solution. A partition-based index is built by first selecting $n_{\text{list}}$ partition representative vectors, typically using $k$-means on a training subset $P_{\text{train}}\subset P$. Each data point is then assigned to the nearest representative, forming disjoint partitions. These assignments are typically stored in an inverted file (IVF) structure (Sivic and Zisserman, 2003). A widely adopted implementation of IVF indexes is the FAISS library from Meta (Douze et al., 2024), which we use as a baseline. See Section 6 for other ANN index types.
3. Solution Overview
We focus on Partition-based IVF indexes because they offer several advantages for deployment in embedding data lakes. They offer the best trade-off in index build time, query response time, and index size (Research, 2023). They have the lowest index build times relative to query throughput (QPS). All of the above are crucial for embedding data lakes, where potentially millions of indexes of various sizes and levels of utilization will have to be constructed and stored.
3.1. Cracking in high dimensions
A well known approach to avoid upfront cost for indexing or organizing data is cracking. In database cracking (Idreos et al., 2007), there is a final state where a single column is sorted. Over time, fewer points are moved, and over shorter distances, reducing the overhead of the cracking. Additionally, in a naive implementation, the physical reorganization happens with every incoming query.
In high-dimensional spaces, the assumption of a fully sorted state no longer holds, and space filling curves, like Z-order (Samet, 1984), do not scale past a few dimensions. Additionally, enforcing reorganization after every query becomes increasingly difficult, as it is a more expensive operation. Moving points between inverted lists incurs random access overhead, with no eventual locality benefits. High-dimensional vectors require moving multiple bytes per dimension, increasing data movement costs. Furthermore, inverted lists, are often implemented as memory-aligned contiguous arrays to allow for efficient SIMD-based kernels, and they may require linear-time compaction even for a single vector to move. These factors contribute to the significant latency overhead, making physical reorganization after every query difficult.
Based on this, the challenge to address is as follows A cracking-based IVF index must efficiently manage the significantly higher overhead of physical reorganization in high dimensions, while minimizing time-to-query, continuously improving the index, and ensuring minimum cumulative time across all scales.
3.2. Outline of the approach
As Figure 2(a) illustrates, index construction can be divided into two operations that independently improve index performance: (i) increasing the number of partitions by randomly selecting partition representatives and (ii) refining these representatives into well-placed centroids using k-means training, which helps to distribute points more evenly across partitions and improve the overall representativeness. Although both improve the index independently, the best results consistently come after reaching a maximum index size, even with randomly initialized representatives, followed by k-means training to optimize their placement. Thus, the optimal approach is to first increase the number of partitions up to a limit and then refine their locations using k-means.
Observation 1 behind the approach: Index construction can be decoupled into two distinct build operations, which can improve the index independently and be separated in time. .
Existing partition-based IVF vector search indexes apply build operations evenly across the vector space, with centroids sampled uniformly and K-means refinements performed simultaneously throughout. While simple and intuitive, this approach assumes equal importance across the space, and requires that both operations are applied at the same time on the entire space leading to larger than necessary index build times. However, vector search workloads can exhibit highly skewed access patterns, where certain regions are significantly more important than others. For instance, in an industrial workload (Mohoney et al., 2023, 2024), up to 85% of centroids are never accessed during search, which would make partitioning or training efforts in those regions unnecessary. As we show in Figure 2(b), skewed query distributions also appear in well known open-source vector search datasets (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016). Even on open source datasets showcasing moderate amounts of skew, where $80\%$ of the total queries access $60\%$ of the partitions, the disparity between the most frequently and least frequently accessed partitions is upwards of 50 times. There is a fundamental inefficiency in current uniform indexing approaches: a large portion of the vector space may rarely accessed, yet the indexing effort is evenly distributed across the space, and always paid upfront.
Observation 2 behind the approach: Just as index construction can be separated in time, it can also be separated in space. The index build operations do not have to be applied uniformly across the vector space; instead, they can be focused on regions that are accessed more frequently. Since certain areas of the vector space are more critical than others, adapting index construction to follow the query distribution is a natural strategy to optimize efficiency and accelerate indexing.
Lastly, in partition-based IVF indexes, a query first identifies the nearest partitions and then scans all points within those partitions to retrieve the $k$ nearest neighbors. This involves the same distance computations needed to update the index, which naturally leads to a cracking-based approach, where index refinement occurs as a side effect of query execution.
Observation 3 behind the approach: We can further amortize the cost of specific build operations, such as assigning vectors to partitions and identifying the local regions to crack and refine, Figure 2(c). We can reuse computations from search operations to perform index construction as a side effect of query execution.
3.3. Outline of the solution
Based on these insights, we propose a partition-based crackable IVF index for ANN search. It is designed to remain near the optimal minimum cumulative time across the number of queries submitted (Figure 1), improving as the workload increases, and addressing the core issue of indexing parts of the search space that are not used.
CrackIVF employs two localized and independent build operations: CRACK and REFINE. The former introduces new partitions, while the latter applies a localized K-means to improve them. These operations run independently across both time and vector space.
We directly follow the query distribution, where queries themselves act as candidates for CRACK and REFINE, determining where new partitions should be introduced, so they are not randomly sampled, and which local regions need refinement. Intermediate results from search operations, such as the distances computed to nearby points, can be re-used to amortize the cost of assignments. Using the queries to create partitions, rather than selecting random ones, allows distances computed during search to reused, steal points from the local region, that are closer to the query than their current assignment. Additionally, using the queries as candidates for CRACK and REFINE enables CrackIVF to exactly follow the query distribution in terms of when and where to execute build operations, as partitions that receive a larger number of queries will proportionally be candidates for cracking and refinement by the same amount.
Finally, to mitigate the high overhead of storage reorganization in high-dimensional IVF indexes, we do not perform build operations after every query. Instead, CRACK operations are buffered and committed when beneficial, while REFINE is executed eagerly but infrequently to improve local regions that queries are visiting when necessary. To automate these decisions, we introduce two control mechanisms. The first is a set of heuristic rules, that determine where in the vector space cracks should be buffered or refinements executed. The second is a cost-based approach that estimates the latency overhead of a build operation and constrains it relative to the cumulative runtime of the index since initialization. As more queries arrive and the index is used more frequently, the budget for build operations increases proportionally, allowing the index to grow and refine in sync with the query workload.
Symbol Description
$|X|$ Cardinality, e.g., $|X|=N$ if $X\in\mathbb{R}^{N\times d}$
[…]
$P\in\mathbb{R}^{|P|\times d}$ Data points, $d$-dimensional
$C\in\mathbb{R}^{|C|\times d}$ Committed cracks
$C_{\text{local}}\in\mathbb{Z}^{nprobe}$ Local region centroid IDs
$P_{\text{train}}$ K-means training points, where $|P_{\text{train}}|=\min(|P_{\text{local}}|,|C_{\text{local}}|\times\max_{\text{points}})$
$n_{\text{iter}}$ K-means iterations
$\max_{\text{points}}$ K-means max training sample per centroid
$nlist$ Total inverted lists storing $P$, where $nlist=|C|$
$Q\in\mathbb{R}^{b_{s}\times d}$ Query set per search
$D\in\mathbb{R}^{b_{s}\times|r|},\,I\in\mathbb{Z}^{b_{s}\times|r|}$ kNN distances and point indices from SEARCH, where $|r|\in[k,|P_{\text{visited}}|]$
$bs$ Total queries batched in single SEARCH
$nprobe$ Number of partitions scanned per SEARCH
$k$ Number of nearest neighbors to return
$A_{P}^{\ast}\in\mathbb{Z}^{|P|}$ Point assignments to centroid IDs
$D_{P}^{\ast}\in\mathbb{R}^{|P|}$ Point distances to centroids
$H_{C}^{\ast}\in\mathbb{N}^{|C|}$ Histogram of points per centroid
[…]
Table 1. Notation used in algorithms and cost model.
4. Crack-IVF
CrackIVF, is a partition-based crackable IVF index implemented in FAISS (Douze et al., 2024). The core logic of the algorithm and operations are covered in Section 4.1 and Algorithm 1. Control mechanisms, including heuristic rules and cost models, are described in Section 4.2 and Section 4.3. They aid in making accurate decisions about where and when to perform the build operations, CRACK and REFINE, which are shown in Algorithm 2 and Algorithm 3. The notation used in this section is summarized in Table 1.
4.1. Algorithm and Operations
CrackIVF performs index construction and storage reorganization operations as a side-effect of query execution to incrementally improve index performance. It has three core operations, $SEARCH$ finds the k-NN results to return to the user, and $CRACK$, $REFINE$ are the build operations that perform physical reorganization of the IVF structure as well as change the index state.
SEARCH: The search procedure is common among partition based IVF indexes, where the local region to the query is identified by the nearest partitions, whose point assignments are subsequently scanned to find the final $k$ nearest neighbors of the query. In our case, the result are point ids, $I$, and distances $D$ to the query, of the visited points in the local region sorted by similarity, together with the ids of the cracks that the query visited, $C_{visited}$, as seen in Line 7 of Algorithm 1.
Algorithm 1 SearchAndCrack
0:  $k,Q,\alpha,State_{dyn},State_{true},T_{buld},T_{search}$
1:  \\ Step 1: SEARCH
2:  $T_{\text{start\_search}}\leftarrow\textsc{CurrentTime}()$
3:  $(D,I,C_{visited})\leftarrow\text{{SEARCH}}(\textit{Q},k)$
4:  $T_{\text{search}}\leftarrow T_{\text{search}}+(\textsc{CurrentTime}()-T_{\text{start\_search}})$
5:  \\ Step 2: CRACK Decision (Where to crack, see §4.2)
6:  for each query $q\in\textit{Q}$ do
7:     $(D_{\text{local}},I_{\text{local}},C_{\text{local}})\leftarrow(D[q],I[q],C_{visited}[q])$
8:     $I_{\text{steal}}\leftarrow\{p\in I_{\text{local}}\mid D_{\text{local}}(p)\mathrel{\text{op}}D_{P}^{\text{dyn}}(p)\}$op = $<$ for L2, $>$ for IP}
9:     $good\_crack\_candidate$ = CrackHeuristic($I_{steal},\text{State}_{\text{dyn}}$)
10:     if $good\_crack\_candidate$ then
11:        $C_{\text{buffered}}\leftarrow C_{\text{buffered}}\cup\{q\}$
12:        $I_{\text{buffered}}\leftarrow I_{\text{buffered}}\cup I_{\text{steal}}$
13:        \\ update dynamic state
14:     end if
15:  end for
16:  \\ Step 3: CRACK Decision (When to Commit, see §4.3)
[…]
18:  $can\_afford\_crack\leftarrow(T_{\text{build}}+\hat{T}_{\text{CRACK}}\leq\alpha\cdot(T_{\text{build}}+T_{\text{search}}+\hat{T}_{\text{CRACK}}))$
19:  $enough\_buffered\leftarrow(|C_{\text{buffered}}|>C_{\min})$
20:  if $can\_afford\_crack$ and $enough\_buffered$ then
21:     \\ UNDO if REFINE invalidated previously good cracks
22:     $C_{\text{buffered}}\leftarrow\{c\in C_{\text{buffered}}\mid\textsc{CrackHeuristic}(c)\}$
23:     $\forall p\text{ where }A_{P}^{\text{dyn}}(p)\notin C_{\text{buffered}},A_{P}^{\text{dyn}}(p)\leftarrow A_{P}^{\text{true}}(p)$
24:     $I_{\text{buffered}}\leftarrow\{p\mid A_{P}^{\text{dyn}}(p)\neq A_{P}^{\text{true}}(p)\}$
25:     $T_{\text{start\_crack}}\leftarrow\textsc{CurrentTime}()$
26:     CRACK$(C_{\text{buffered}},I_{\text{buffered}})$Algorithm 2}
27:     $T_{\text{build}}\leftarrow T_{\text{build}}+(\textsc{CurrentTime}()-T_{\text{start\_crack}})$
28:     \\ dynamic index state now matches true state
29:  else
30:     \\ Step 4: REFINE Decision (Where & When, see §4.2 and §4.3)
31:     for each query $q\in\textit{Q}$ do
[…]
33:        $can\_afford\_refine\leftarrow(T_{\text{build}}+\hat{T}_{\text{REFINE}}\leq\alpha\cdot(T_{\text{build}}+T_{\text{search}}+\hat{T}_{\text{REFINE}}))$
34:        if $can\_afford\_refine$ then
35:           $good\_refine\_candidate\leftarrow\textsc{RefineHeuristic}(C_{\text{local}},\text{State}_{\text{true}})$
36:           if $good\_refine\_candidate$ then
37:              $T_{\text{start\_refine}}\leftarrow\textsc{CurrentTime}()$Start timing REFINE}
38:              REFINE()($\dots$)( … )
39:              $T_{\text{build}}\leftarrow T_{\text{build}}+(\textsc{CurrentTime}()-T_{\text{start\_refine}})$
40:              \\ update dynamic state and true state
41:           end if
42:        end if
43:     end for
44:  end if
45:  return  $D[:,:k],I[:,:k]$
CRACK: Physical reorganization of points between inverted lists, and the mutation of the index does not happen after each query. Each incoming query represents a crack candidate, which is evaluated by heuristic rules to classify it as good or bad (Section 4.2) (Line 9 Algorithm 1). Only good cracks are further buffered and potentially committed at a later time. Each crack is essentially a new partition, where the points assigned to the partition have been “stolen” based on their proximity to the query. We reuse the results of the SEARCH operations to compare if the distance of a point to the query is closer than the distance of the point to its current assignment, which could be another buffered crack. If it is, then it’s beneficial to reassign the point to the new crack, and this assignment is also buffered, to be committed when $CRACK$ is executed, as explained in Section 4.3. $CRACK$ is the procedure performing the physical reorganization, syncing the buffered state of the index with the currently working state. The core individual kernels of the procedure are sequentially executed (Algorithm 2). It adds a new inverted list for each buffered crack, moves stolen points to their new assignments, and updates the partition representatives to be the centroid of the points based on their new assignments. This updates the location of the crack representatives, whose point assignments were stolen, in order for their new location to match the underlying assignment distribution. This procedure is only executed when enough budget for build operations has been accumulated through the usage of the index (Line 18 Algorithm 1). This control mechanism ensures that CrackIVF does not spend a disproportionate amount of time on CRACK operations. The parameter $\alpha$, the ratio of total build operations time to total search time, is what controls the available budget. Thus, since $CRACK$ is lazy, the heuristics governing where a crack is added and the cost-based control mechanism for when it is added are independent.
Algorithm 2 CRACK
0:  $C_{\text{buffered}}$, $I_{\text{buffered}},State_{dyn},State_{true}$
0:  New cracks committed; points reassigned; index state updated.
1:  Get Local Region: Retrieve all points indexed by $I_{\text{buffered}}$, i.e., all data points that cracks in $C_{\text{buffered}}$ visited.
2:  Commit Reorg (w/ $\mathbf{A_{P}^{\text{dyn}}}$) : Reorganize storage. Commit $C_{\text{buffered}}$ to the index. Reassign points to inverted lists using tracked assignments $A_{P}^{\text{dyn}}$
3:  Update Centroids: Recompute each crack’s representative to be the centroid of the points assigned to it.
4:  Update Index State: Synchronize $State_{dyn}$ and $State_{true}$ to reflect the changes.
REFINE: Refinement is the second operation to improve the index. It is performed eagerly, focusing on the local region visited by the query. The goal is to immediately correct suboptimal regions in the current index. Thus, the buffered states and buffered cracks within the local region are not considered. The decision to refine a region is based on heuristic rules that evaluate the imbalance of point assignments in local cracks. Additionally, executing a refinement operation depends on an estimate of the runtime cost, determined using a cost model specific to $REFINE$ (Line 33 Algorithm 1).
Unlike cracking, the opportunity to refine a local region is transient. The heuristic rules must decide that the currently visited local region is a good candidate for refinement, and at the same time, there must be sufficient computational budget to execute the operation. If where and when to execute a refine do not simultaneously agree, the next opportunity for the same region will arise only when another query accesses it.
$REFINE$, as described in Algorithm 3, executes a sequence of kernels implementing a local K-means variation and performs physical reorganization. This process improves the placement of crack representatives within the local region and reassigns points to their nearest representative. Unlike $CRACK$, $REFINE$ incurs additional computational costs as it does not rely on precomputed buffered assignments. Instead, assignments are determined dynamically upon completion of the local K-means process.
Algorithm 3 REFINE
0:  $C_{local},State_{dyn},State_{true}$
0:  Refinement local region; points reassigned; index state updated.
1:  Get Local Region: Retrieve all points $P_{local}$ in local region.
2:  Local K-Means: Refine region of $C_{\text{local}}$, producing $C^{\prime}_{\text{local}}$ centroids.
3:  Commit Reorg (w/o $\mathbf{A_{P}^{\text{dyn}}}$): Replacing $C_{\text{local}}$ with refined $C^{\prime}_{\text{local}}$. Compute new local point assignments. Reorganize storage, with local reassignment of $P_{local}$ to inverted lists.
4:  Update Index State: Synchronize $State_{dyn}$ and $State_{true}$ to reflect the changes.
Index State: Since CRACK is a lazy operation, the index can exist in two concurrent states, the true state and a dynamic state. The true state reflects the current structure of the index, while the dynamic one tracks its expected future structure for when $CRACK$ is executed, which synchronizes the two states. Maintaining both is essential. $REFINE$ operations directly alter the current index state, while cracking decisions are made based on the dynamic one, leading to scenarios that require us to undo buffered crack decisions. Initially, the heuristic rules considered some crack candidates as beneficial, and CrackIVF buffered them, but they have since been altered. Rolling back the dynamic state avoids bad cracks from being committed and having a situation where we commit decisions to the index that we know to be suboptimal (Line 21 Algorithm 1).
To track the dynamic, uncommitted, state of the index, we define:
(1) $\text{State}_{\text{dyn}}=(A_{P}^{\text{dyn}},D_{P}^{\text{dyn}},H_{C}^{\text{dyn}})$
which consists of dynamic assignments $A_{P}^{\text{dyn}}$, dynamic point distances $D_{P}^{\text{dyn}}$, and a histogram $H_{C}^{\text{dyn}}$ of true and buffered crack sizes.
Similarly, the true state is defined as:
(2) $\text{State}_{\text{true}}=(A_{P}^{\text{true}},D_{P}^{\text{true}},H_{C}^{\text{true}})$
representing the true, committed state of assignments, distances, and crack sizes in the index.
The initialization of the above states happens during the index setup time. More specifically, in our FAISS-based implementation, during an “.add()“ operation. This is when the data vectors are assigned to their nearest partitions by calculating the distances to each partition representative. We extend this operation to return to the assignments and distances computed, which were an intermediate result. The state held in $A_{P}^{\text{dyn}}$,$D_{P}^{\text{dyn}}$ and $H_{C}^{\text{dyn}}$, is only consistent with the actual state of the index after $CRACK$ operations and before buffering any cracks. If no cracks are buffered and only $REFINE$ operations are executed, the dynamic and true state are always consistent. All of the above state metadata is used after the SEARCH operation, along with the search results $I,D,C_{local}$ to classify each query as a good or bad crack candidate (Section 4.2).
Rollback of buffered state: Consider the following scenario. A $CRACK$ operation was just executed so at $t_{0}$ $State_{true}=State_{dyn}$. After a few queries, at $t_{1}$, assume that 2 good crack candidates have been buffered, both in the $C_{local}$ region of the vector space and thus now $State_{true}\neq State_{dyn}$. At a later point, $t_{2}$, a $REFINE$ is executed, also in $C_{local}$ refining it to $C^{\prime}_{local}$ and changing $S_{true}$. Since $REFINE$ operates to improve the current state eagerly, the decision that the buffered cracks are good was taken prior, at $t_{1}$ based on $C_{local}$, and thus they may now have had their points stolen back, and assigned to the refined cracks in $C^{\prime}_{local}$. Thus even though still $|C_{buffered}|=2$, it may now contain bad cracks. So when at $t_{3}$, it is a good time to execute a $CRACK$ operation again, all buffered cracks are re-evaluated, and only the good cracks that remain are committed, ensuring that in the end, we are back to $State_{true}=State_{dyn}$. This logic can be seen in lines 21-25 of Algorithm 1.
Memory Overhead and Optimizations: The additional memory cost of maintaining $State_{\text{dyn}}$ and $State_{\text{true}}$ is $O(4|P|+2|C|)$, which is negligible since $d\gg 4$, so in our in-memory implementation that data alone are $O(|P|*d)$. Additionally, these array structures are freed once the index converges or incoming queries stop. In our case, it enables an efficient compute-storage trade-off, deferring cracking costs instead of paying them per query.
There are a number of optimizations that a fully fledged implementation can make use of. For embedding lakes, where vectors are stored on disk, state tracking arrays can be stored alongside data points instead of in the index (irrelevant for our in-memory FAISS implementation). Since queries for an individual index are typically infrequent, $State_{\text{dyn}}$ can be committed or flushed between queries, while $State_{\text{true}}$ is recomputed as needed. Full memory overhead is only necessary during bursty query loads, where cracking must be buffered to reduce cumulative overhead and prevent long tail response times. Finally, the cracking overhead can be hidden from query response times since results are available before any cracking operations begin. A separate thread pool can handle cracking operations while queries use the current index state, provided $CRACK$ and $REFINE$ remain atomic to prevent inconsistencies. Prior work on the related problem of index maintenance has demonstrated similar approaches to minimizing the overhead of mutating the ANN index state (Xu et al., 2023) with concurrent operations, we further cover these approaches in Section 6.
Our work assumes a continuous query load at the system’s capacity, leveraging inter-query parallelism. As a result, the memory overhead for index state tracking is unavoidable in our implementation if we aim to maintain high performance.
4.2. Where to apply build operations?
There is a constant stream of crack and refinement candidates, and as observed in Section 3, not all regions of the vector space are equally important, and not all operations applied lead to improvement of index performance. Thus, a set of rules is needed to distinguish good candidates from bad ones. The heuristic rules provided in this section aim to distill the learning from our initial observations to capture independent cases where we want to add a new crack or refine a region. At this stage, they are not an exhaustive list, as we plan to extend them in future work.
Where to CRACK: The CrackHeuristic determines whether an incoming query is a good or bad candidate for cracking, as shown in Algorithm 1. It evaluates the stolen point IDs $I_{\text{steal}}$ and the dynamic state $State_{\text{dyn}}$ to return a True/False decision. If classified as good, the query is buffered for a future $CRACK$.
By default, we aim to expand the index size, as our insights suggest that optimal performance is typically achieved at some maximum number of partitions. Since $REFINE$ does not add partitions, we assume every incoming query is a good candidate to have a constant stream of candidates that can expand the size of a CrackIVF index and only reject them if they pass any of our heuristic rules.
CrackHeuristic: A query is considered a bad candidate (returning False) if any of the following conditions hold:
Don’t Crack: Rule 1 rejects queries with too few stolen points:
[…]
Don’t Crack: Rule 2 prevents excessive partitioning, constraining the number of partitions within a region relative to the number of points:
[…]
Where to REFINE: The RefineHeuristic determines whether a local region in the index is a good candidate for refinement, returning a $True/False$ value for if a refinement should take place. Unlike $CRACK$, where the latency overhead is amortized by buffering, $REFINE$ is applied eagerly and is also an expensive operation, running K-Means locally, calculating new point assignments, and reorganizing the physical storage. Additionally, if the index continues to grow in the future, any previously refined region may need further refinement, making premature refinements inefficient. It is much better to relax rules for when to refine when the index has reached its maximum size and has also accumulated a larger cost budget.
Therefore, we do not want to be calling $REFINE$ repeatedly, unless necessary, and our default behavior is to assume that every incoming query local region is NOT a good candidate for refinement unless it satisfies specific heuristic rules that indicate a significant local partition imbalance. Imbalance affects tail response time negatively and increases the variance since even though queries may be accessing the same number of partitions, they may be scanning a vastly different number of points. Additionally, running K-means improves the centroid location of the crack representatives, making them better fit the underlying data distribution, which also has downstream benefits (Mohoney et al., 2024).
Various variance measures exist for detecting imbalances in cluster assignments (Hastie et al., 2009). We selected the following subset as they capture different aspects of imbalance, both locally and globally.
RefineHeuristic: A local region is considered a good candidate for refinement (returning True) if any of the following conditions hold:
Refine: Rule 1 captures local imbalance using the coefficient of variation:
[…]
where $\sigma_{\text{local}}$ and $\mu_{\text{local}}$ are the standard deviation and mean of cluster sizes in the local region. A high coefficient of variation indicates substantial local imbalance, warranting refinement.
Refine: Rule 2 captures extreme values using local spread:
[…]
This ensures that refinement is triggered when large disparities exist between the smallest and largest clusters within a local region.
Refine: Rule 3 captures imbalance based on the global distribution:
[…]
where $c_{1},c_{2}$ are cracks in the local region, and $S(c)=H_{C}^{\text{true}}[c]$ represents the size (i.e., number of assignments) of crack $c$.
[…]
These thresholds ensure that refinement is applied only when clear imbalance exists, maintaining index efficiency. An optimized implementation could relax these rules when more computational budget is available, allowing refinement of less imbalanced regions.
4.3. When to apply build operations?
To mitigate the overhead of constant physical reorganization, which is expensive in high-dimensional IVF indexes, we use a build budget that restricts the fraction of time dedicated to $CRACK$ and $REFINE$ operations. We define $T_{\text{build}}$ as the total measured time spent on previous build operations, which include historical $CRACK$ and $REFINE$. Similarly, $T_{\text{search}}$ represents the cumulative time spent on all past $SEARCH$ operations. When considering a potential build operation, we estimate its execution cost using $\hat{T}_{f}$, which captures the expected time required to perform either a $CRACK$ operation under the current dynamic state ($State_{dyn}$) or a $REFINE$ operation on the local region of the current refine candidate. Here, $f\in\{\textsc{CRACK},\textsc{REFINE}\}$ specifies the build operation considered.
To enforce the budget constraint, we ensure that the total time spent on past and projected build operations remains within a fraction $\alpha$ of the system’s total runtime:
(8) $T_{\text{build}}+\hat{T}_{f}\leq\alpha\cdot(T_{\text{build}}+T_{\text{search}}+\hat{T}_{f}).$
$T_{\text{build}}$ and $T_{\text{search}}$ are directly measured from execution, while $\hat{T}_{f}$ is derived from a predictive cost model, which is explained in the following subsection. The parameter $\alpha$ governs the proportion of total execution time that can be allocated to indexing operations, ensuring a balanced trade-off between indexing efficiency and query performance. This approach allows for a consistent and adaptive mechanism to regulate both cracking and refining decisions under varying workload conditions.
Varying $\alpha$ is a way of expressing an expectation about the number of queries that will arrive. A large $\alpha$ allocates a larger build budget, which is useful if we are optimistic about receiving a large number of queries in the future, thus spending more time to build an index is beneficial. A smaller $\alpha$ restricts the build budget, indicating that we are pessimistic and do not expect many more queries to arrive. To have a balance, in all our experiments, we set $\alpha=0.5$, to ensure a maximum of 50% of the total time at any given point is spent on build operations. This budget ensures that the build operation overhead does not dominate the overall execution time. As more queries arrive, CrackIVF will progressively have a larger budget to apply for physical reorganization operations. In this way, we can remain near the optimal minimum of cumulative time across all query scales. Finally, there is a minimum size constraint on the number of buffered cracks before a $CRACK$ is executed. Specifically, in our implementation, we set that minimum to 20% of the current index size, $|C_{buffered}|>C_{min}$, where $C_{min}=0.2*|C|$ (Line 19 Algorithm 1). This helps avoid situations where a $REFINE$ operation applied on the current index achieves similar performance improvements, while allowing us to keeping the current index size without increasing it.
Predictive Cost Model
The cost estimation of both CRACK and REFINE operations happens after every incoming query. Thus, it is crucial to maintain low latency when making a prediction while at the same time maintaining an accurate enough estimate to satisfy our usage needs. To achieve this balance, we employ a simple first-order linear model and fit it using multivariate regression (Jain, 1990).
Each build function $f$ comprises a sequence of procedures, or kernels, indexed by $i$ and executed sequentially, as shown in Algorithm 2 and Algorithm 3. Among these, updating the index state is negligible, and the following kernels are the dominant factors of each function’s execution time:
CRACK: (i) Get Local Region, (ii) Commit Reorg (w/ $A_{P}^{\text{dyn}}$), (iii) Update Centroids.
REFINE: (i) Get Local Region, (ii) Local k-Means, (iii) Commit Reorg (w/o $A_{P}^{\text{dyn}}$).
Since the above kernels in our implementation execute sequentially without overlap, the total execution time $T_{f}$ is simply the sum of the execution times of its kernels.
(9) $T_{f}=\sum_{i}T_{f,i},\quad\forall f\in\{\textit{CRACK},\textit{REFINE}\}.$
Each kernel $i$ contributes a latency $T_{f,i}$, which we model as a linear function of its two dominant cost factors: computational complexity and data movement. Specifically, for each kernel $i$ in function $f$, we train the following simple predictive model:
(10) $T_{f,i}=w_{f,i}^{(1)}\cdot C_{f,i}+w_{f,i}^{(2)}\cdot D_{f,i}+b_{f,i},$
where $C_{f,i}$ represents the dominant computational complexity of kernel $i$, and $D_{f,i}$ denotes the dominant data movement cost. The terms $w_{f,i}^{(1)}$ and $w_{f,i}^{(2)}$ are learned coefficients corresponding to computation and data movement, respectively. Finally, $b_{f,i}$ is a learned parameter that accounts for fixed kernel execution overheads.
We gather measurement with micro-benchmarks for each kernel, by varying input parameters and measurig execution latency to fit a regression model, learning $w_{f,i}^{(1)},w_{f,i}^{(2)},b_{f,i}$.
Our approach eliminates the need for handcrafted analytical performance models that map the kernels into the number of cycles per operation and are tied to specific implementations and hardware configurations, making them hard to adapt or extend. With our method, switching hardware or optimizing kernels only requires re-running microbenchmarks and refitting the linear model. Another advantage is the low computational cost of inference, enabling continuous use during query execution. However, modeling execution time as a simple linear function has limitations. While it captures dominant compute and memory trends, it overlooks system-level nonlinearities. Factors like caching effects, NUMA constraints, SIMD optimizations (e.g., in FAISS), and thread contention introduce discrete performance variations that a linear model cannot fully express.
Kernel Compute Data Movement
Get Local Region* $|C_{\text{buffered}}|\times|C|$ $|P_{\text{local}}|\times d$
Commit Reorg (w/o $A_{P}^{\text{dyn}}$) $|P_{\text{local}}|\times|C_{\text{local}}|\times d+|C|$ $|P_{\text{local}}|\times d$
Commit Reorg (w/ $A_{P}^{dyn}$) $|C|$ $|P_{\text{local}}|\times d$
Update Centroids $|P|\times d$ $(|P|+|C|)\times d$
Local K-Means $n_{\text{iter}}\times|P_{train}|\times|C_{\text{local}}|\times d$ $|C|\times d$
Table 2. Features we used for each kernel. * For REFINE, $C_{\text{buffered}}=1$
Nevertheless, our empirical results show that the cost model performs well across all benchmarks. Importantly, we always use the true measured times $T_{\text{build}}$ and $T_{\text{search}}$ in our budget constraint Equation 8. This ensuring any overestimation by the predictive model is corrected after execution, and an underestimation will hold any future build operations until sufficient queries arrive, increasing the total budget. In our testing, the model achieved moderate to high $R^{2}$, explaining between 40 and 95% of the total variation, with the $RMSE$ remaining in the millisecond to low second range. We find this to be acceptable performance from our model since we benefit from low inference latency, and the cost budget-based control mechanism works effectively on all datasets, as we show in our experiments.
5. Experiments
5.1. Experimental Setup
We run on a dual-socket AMD EPYC 7V13 system (128 cores, 3.7 GHz, 512GB RAM) from the AMD-ETHZ HACC cluster (Moya et al., 2023a, b, 2024). Unless otherwise stated, experiments use 16 cores, the number at which the memory bandwidth saturates, and use an equal query batch size to benefit from inter-query parallelism.
Table 3. AV-Tree vs. CrackIVF
Cumulative Time at i-th Query (seconds)
$10^{0}$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ (Total)
AV-Tree 0.078 0.747 7.601 75.369 825.709 1264.175
CrackIVF 1.107 1.166 1.738 15.983 98.295 516.881
Datasets: We evaluate on standard ANN search benchmarks for text and image-based retrieval, GloVe(Pennington et al., 2014), SIFT(Jégou et al., 2011), DEEP(Babenko and Lempitsky, 2016), and Last.fm(Bertin-Mahieux et al., 2011). The latter, being an example with a high skew, is used for music recommendation. The SIFT1M and SIFT10M, contain 128-dimensional SIFT descriptors, while DEEP10M consists of 96-dimensional embeddings from a fully connected GoogLeNet model (Szegedy et al., 2015). These datasets are subsets of the original billion-scale collections, with 1M and 10M referring to the number of data points in each slice. Both use L2 distance as a metric. GloVe, sourced from ANN-Benchmarks (Aumüller et al., 2020), represents text retrieval. It contains 1.18M varying dimensional data points and uses cosine similarity, which we implement by l2-normalizing and computing inner product (IP) similarity. Last.fm is also taken from (Aumüller et al., 2020); it contains 64-dimensional vectors and uses the IP similarity metric. We found that the provided query set is highly skewed, allowing us to test how CrackIVF behaves in such a scenario. Last.fm provides 50k unique queries, while the other datasets provide 10K unique queries. Thus, we replicate them until we match our target scales (up to 1M).
Baselines: Our comparisons cover both exact nearest neighbor (ENN) and approximate nearest neighbor (ANN) search approaches. For ENN, we use Brute Force, i.e., a linear scan of the data and AV-Tree (Lampropoulos et al., 2023), as a cracking baseline. For ANN search, we compare against FAISS IVF indexes, which in our figures are indicated as IVFFlat, with the number of partitions chosen in parentheses, e.g., IVFFlat (1000). We vary the number of partitions from 1,000 to 16,000, following FAISS guidelines (Facebook Research, 2025), which recommend this range for our dataset scales. Our method, CrackIVF is initialized with 100 partitions to maintain near brute force startup time.
All experiments have the same fixed parameters, with values set as explained in (Sections 4.2, 4.3). The $\alpha$ parameter to control the budget remains fixed to $0.5$. Allowing, a build budget equal to 50% of the total time. Finally, for the partition-based indices, we have CrackIVF and IVFFlat baselines, where the $nprobe$ is set to always achieve 90–95% Recall@10. The QPS-Recall plots show the Pareto frontier when varying $nprobe$ across all indexes.
(a) SIFT1M: QPS vs. Recall
(b) SIFT1M: Time per Query
(c) SIFT10M: QPS vs. Recall
(d) SIFT10M: Time per Query
(e) Last.fm: QPS vs. Recall
(f) Last.fm: Time per Query
(g) Glove50: QPS vs. Recall
(h) Glove50: Time per Query
(i) Glove100: QPS vs. Recall
(j) Glove100: Time per Query
(k) Deep10M: QPS vs. Recall
(l) Deep10M: Time per Query
Figure 3. Queries Per Second (QPS) vs. Recall and Time per each Query batch for the entire trace across different datasets.
5.2. Comparison with AV-Tree:
To compare with AV-Tree (Lampropoulos et al., 2023) fairly, we set our CrackIVF to run with a single thread and make sure we probe enough partitions to get an average of more than $99\%$ recall across all query scales. We set the cracking threshold of AV-Tree to 128, which we find gives the best total runtime. We show the cumulative time for a trace of $100$K queries over the SIFT 1M dataset in Table 3. Unlike ours, AV-Tree does not pre-build the index, therefore, it has an advantage over us for the first 10 queries. However, our method processes each query faster and outperforms AV-Tree by a large margin after the first few queries. The cumulative time of AV-Tree increases at a slower rate with more queries, meaning that the query time is getting faster, and AV-Tree is improving over time. Nevertheless, our CrackIVF is 2.45x faster than AV-Tree in the end, even on a single-threaded ENN search. This shows the effectiveness of ANN approaches, where a less than 0.01% Recall error leads to a large gain in performance. For the rest of the experiments, we focus on the ANN search setting.
5.3. Does CrackIVF improve over time?
To measure how CrackIVF improves over time, we provide two graphs for each of the datasets and group them in Figure 3. The QPS vs. Recall plots measure the maximum achievable Queries Per Second (QPS) across varying Recall targets. This is the trade-off under which the ANN search operates. Having a large gain in search speed (QPS) for a small decrease in search quality (Recall). By the end of the 1 million query trace, indicated by CrackIVF-END (orange line), we are consistently at the Pareto frontier of QPS-Recall across on all datasets. On SIFT10M and Last.fm, CrackIVF is by far the best-performing index. Specifically on Last.fm, CrackIVF with 1255 final partitions, achieves 50% higher QPS for Recall@10 = 0.9, than a 10x larger IVFFlat(16000) index. We attribute this to CrackIVF’s approach of only allocating new cracks based on the query distribution, and since Last.fm is the dataset with the highest skew, where only  30% of the vector spaces is accessed (Figure 2(b)), we can fully partition that  30% of the space with a relatively small number of cracks to achieve the best performance.
In Figure 3, we also plot the Response Time per query batch across the entire 1 million query trace. We can see that CrackIVF consistently improves over time, showing drops in response time after incremental reorganization operations. The outliers can all be attributed to these $CRACK$ and $REFINE$ reorganization operations. This is evidence for having control mechanisms to ensure expensive operations are only applied when necessary and why CrackIVF relies on buffering crack operations. The downstream of buffering and committing a large number of multiple cracks at once is only visible in one dataset, Last.dm, where it creates a step-like response in the data, as seen in Figure 3(f). This sudden improvement in performance when adding a lot of cracks can also be seen in Figure 2(a), when going from 1000 to 5000 partitions.
5.4. Does CrackIVF Minimize Cumulative Time?
Our core motivation in this work has been to provide an index that can be deployed across data lakes under scenarios where the number of queries that will arrive for each dataset and utilize the index is unknown. The ideal index for such a scenario should be able to minimize the cumulative time spent on both build and search operations in cases where a dataset may receive a few queries as well as millions.
(a) SIFT1M
(b) SIFT10M
(c) Last.fm
(d) Glove50
(e) Glove100
(f) Deep10M
Figure 4. Cumulative time plots across datasets.
We evaluate this specific aspect of index performance, across all datasets, with the cumulative time experiments, shown in Figure 4. For our measurements, we run a trace of 1 million queries for each dataset. In all baseline methods, we include the upfront build cost as the time before they answer the first query. The smaller IVFFlat (1000) indexes have relatively low startup costs but do not scale well past  10k queries since their cumulative time rises sharply due to having higher query batch response times than the alternative. The larger indexes, beyond 5000 partitions, have a huge upfront build cost, which may pay off but only if a dataset receives ¿1 million queries. The “no index“ approach of the Brute-force method has the minimum initialization time and is thus the best choice when the number of queries is low (¡100).
CrackIVF consistently achieves near-minimum cumulative time across all query scales. It achieves up to 1000x lower initialization time than larger indexes (Figure 4(d)) and can consistently answer 100K to 1M queries before comparable indexes finish building. For example on SIFT10M (Figure 4(b)), the initialization time cost is 10x lower than IndexFlat(5000) and 100x lower than IndexFlat(16000), yet it processes 100K and 1M queries respectively, before they baselines indexes complete construction, while at the same time converging to higher QPS-Recall performance. Similar results can be observed in all other datasets.
The fact that CrackIVF achieves the Pareto optimal result in QPS-Recall by the end of our trace means that, by extrapolation, it will always maintain the minimum cumulative time across all query scales beyond 1M queries.
5.5. Control Mechanisms Ablation Study
(a) WHERE=OFF: QPS vs Recall
(b) WHERE=ON: QPS vs Recall
Figure 5. Comparison of QPS-Recall, showing effect of removing the ”where” to CRACK and REFINE mechanism
(a) Cumulative Time (OFF)
(b) Total Time Breakdown (OFF)
(c) Cumulative Time (ON)
(d) Total Time Breakdown (ON)
Figure 6. Comparison of cumulative time and total time breakdown, showing effect of removing the ”when” to CRACK and REFINE mechanism
CrackIVF uses two control mechanisms for the CRACK and REFINE build operations, as detailed in Section 4.2 and Section 4.3. The first mechanism, controlling where they should be applied, uses a set of heuristic rules. The second mechanism, controlling when they are executed, is a cost-budget-based approach combined with a predictive model to balance search and build times. This section demonstrates the importance of these controls in achieving optimal performance and additionally highlights the effect of changing the parameter default value $min\_pts=2$, which changes the heuristic rule “Don’t Crack: Rule 1“ (Section 4.2).
Turning off the control mechanism for “where“: To illustrate the importance and accuracy of the heuristic rules for where to CRACK and REFINE, we switch them OFF and see the effect (Figure 5). Specifically, when “WHERE=OFF”, and no heuristic rule is used to decide where a crack or refine should happen, the index defaults to trying to apply these operations whenever there is enough budget for them (the “when“ mechanism is still ON). This seemingly random addition of new cracks and local refines leads to a final performance that is no better than the initial state of CrackIVF, before any query has been received (Figure 5(a)). On the other hand, when “WHERE=ON“, and our heuristic rules are applied to make decisions, CrackIVF converges to the Pareto optimal QPS-Recall trade-off performance (Figure 5(b)). In both examples, we send an equal number of queries to the index and in the same order.
Turning off the control mechanism for “when“:
The budgeting mechanism, controlling when build operations are executed, helps balance the time spent on build vs search operations. By default, we use the parameter $a=0.5$, i.e., at most 50% of the total time may be spent on $CRACK$ or $REFINE$. The effect of turning this mechanism “WHEN=OFF“ is shown in Figure 6(a) and Figure 6(b). CrackIVF with the mechanism off defaults to cracking and refining after almost every query, and it only stops when the maximum number of partitions is reached (16,000). So even though the cracks and refines happen in regions that require them, they happen disproportionally often compared to the search overhead, leading to a final time spent on build operations that completely dominates the total time. On the other hand, when the mechanism is enabled (“WHEN=ON“), CrackIVF is able to efficiently balance the search and build costs (Figure 6(b) and Figure 6(c)). These runs are for the same 100k query trace.
Varying heuristic rule parameters:
(a) QPS vs Recall ($min\_pts=2$)
(b) QPS vs Recall ($min\_pts=32$)
Figure 7. Comparison of QPS vs Recall for different $min\_pts$ values.
CrackIVF’s heuristic rules use default parameters (Section 4.2). To avoid parameter tuning, we did not vary these values across experiments. However, testing revealed that the $min\_pts$ parameter can impact performance, dependent on the dataset. $min\_pts$ defines the minimum number of points that must be stolen by a new crack for it to be considered a “good“ and subsequently buffered. On the GloVe-25 dataset, changing $min\_pts$ from 2 to 32 notably improved the QPS-Recall performance of the final index (Figure 7). Originally, we set $min\_pts=2$, a very small value, to ensure we avoid adding empty cracks to the index. We assumed that any imbalances could be fixed by future $REFINE$ operations that redistribute points locally, which works for all datasets except for GloVe-25.
Although we can not make a definite statement, we hypothesize that this can be attributed to the fact that this dataset only has 25-dimensional embeddings. This leads to search times that are very low, and subsequently, our build operations budget grows slower. Thus, there are too many imbalances introduced with $min\_pts=2$, and a smaller budget for future $REFINE$ operations to correct them. By increasing $min\_pts$ to 32, we reduced the total number of local imbalances by avoiding to ever commit excessively small cracks, which in turn reduces the number of $REFINE$ needed to reach a good final performance.
A second hypothesis comes from closer inspection of the Figures, in which the better performing index converges to 10056 partition vs the default 16000. (Figure 7(a), Figure 7(b) ). But, we believe because GloVe-25 queries access the entire vector space evenly (Figure 2(b)) and it is a dataset with (1.18M points), we do not think a larger number of partitions hinders performance.
Nevertheless, we show that parameter values do matter across datasets, so index tuning can become a consideration. But for this work, across all the datasets we tested, outside of GloVe-25, the parameter defaults we have chosen have reliable performance.
Takeaways: Our results show that CrackIVF minimizes the cumulative costs across a large variety of individual datasets tested, but to fully grasp the potential performance benefits in embedding lakes, it’s essential to consider the aggregate across all datasets’ cumulative time savings and as the number of queries varies. For systems managing potentially thousands or millions of datasets, all saving are scaled equivalently, and choosing a separate index for each dataset is near-impossible, so the benefit from using CrackIVF is greater, as there is no one-size-fits all solution from the pre-built indexes.
From our experiments, we make the following suggestion: Brute Force should be used as the default method up until a dataset surpasses 10–100 queries, at which point CrackIVF becomes the optimal choice. Even though CrackIVF can start with one partition to simulate a Brute Force approach, it has a slight memory and runtime overhead compared to it, from the setup time and storage of internal object structures and keeping track of the first few partition representatives. At scale, for embedding lakes with millions of datasets, even this small overhead can accumulate. Therefore, we recommend using CrackIVF once a dataset exceeds 100 queries and is thus likely to receive even more.
6. Related Work
Vector Databases: Vector Data Management Systems (VDBMS) (Wang et al., 2021; Weaviate, 2025; Pinecone, 2025; Zhang et al., 2023; Chen et al., 2024; Guo et al., 2022), are systems that support efficient ANN search, and have gained significant popularity in both research and industry. This coincides with the increase in popularity of Large Language Models (LLMs) (Brown et al., 2020), that use Retrieval-Augmented Generation (RAG) techniques (Lewis et al., 2020; Asai et al., 2023), to extend their capabilities beyond what they have been trained on, in order to answer queries based on external data sources. Vector Database Management System applications extend beyond the above use-case, as they are also used as platforms for semantic search (Salton et al., 1975; Mitra et al., 2018), recommendation systems (Li et al., 2017) and product search in e-commerce (Van Gysel et al., 2016). Given the importance of these operations, many existing data management systems are incorporating vector search capabilities (Corporation, 2025; pgvector contributors, 2025). Many of these systems also support extensions to dense vector ANN search, through specialized methods, allowing for specific query types such as filtered search (Gollapudi et al., 2023), sparse vector search (Formal et al., 2021), and hybrid search (Vespa.ai, 2024), which combines vector search with more traditional information retrieval techniques like BM-25 (Robertson et al., 2009). A recent survey of vector database management systems is found in (Pan et al., 2024).
Vector Search Indexes: Approximate Nearest Neighbor (ANN) search is a well-studied field with decades of research, resulting in a large number of indexes. Broadly, the two dominant categories of ANN vector search indexes are partition-based and graph-based indices. Inverted index (IVF) based methods, which fall under partition based category, include SCANN and SOAR (Guo et al., 2020; Sun et al., 2024), which identify key mathematical properties of ANNS and use them to improve IVF index efficiency, SPANN (Chen et al., 2021), which shows how to leverage disk-based storage for search over billion-scale datasets along with supporting updates (Xu et al., 2023), and FAISS-IVF (Douze et al., 2024), an in-memory implementation utilizing GPUs (Johnson et al., 2019) and quantization techniques like PQ (Sivic and Zisserman, 2003; Jegou et al., 2010; Ge et al., 2013). In this work, we focus on IVF methods, and our cracking based approach could be extended for the above implementations. IVF indexes offer the best trade-off between index build time, search performance, and memory overhead, all of which are critical to balance for potential large scale deployments in embedding data lakes. Hash based methods (Andoni et al., 2015; Indyk and Motwani, 1998; Datar et al., 2004; Jain et al., 2008; Wang et al., 2017) and tree based methods (Muja and Lowe, 2014; Liu et al., 2004; Wang et al., 2013) also broadly fall under the partition-based index category. Finally, graph-based methods (Malkov and Yashunin, 2018; Malkov et al., 2014; Wang et al., 2012; Dong et al., 2011; Wang et al., 2013; Hajebi et al., 2011) currently deliver the state-of-the-art performance in terms of the Queries Per Second vs. Recall trade-off, but this comes at high memory and index construction cost. Thus, graph based indexes are for now ill-suited for scenarios where minimizing the index construction cost and low memory overhead is a important, for example in embedding lakes where a very large number of indexes is expected to be constructed.
Adaptive indexes for specialized data types: Adaptive indexing techniques, such as Dumpy (Wang et al., 2023, 2024) and ADS (Zoumpatianos et al., 2014), have demonstrated significant effectiveness in data series similarity search. These indexes are specifically designed for time-series data and include methods like Symbolic Aggregate Approximation (SAX) (Lin et al., 2003) and indexable SAX (iSAX) (Shieh and Keogh, 2008). In our work, we focus on ANN indexing methods that are the standard used by RAG systems, and due to their broad applicability across various data modalities which lack sequential relationships. That said, specialized, adaptive and incrementally constructed indexes are relevant candidates for embedding lake deployments, particularly for the subset of the datasets with inherent sequential dependencies. This includes energy, weather and financial data, as well as audio and video recordings (Raza et al., 2015; Mirylenka et al., 2016; Shasha, 1999; Kashino et al., 1999).
ANN index maintenance: Work on index maintenance focuses on mitigating performance degradation caused by data distribution shifts due to data updates (insertions and deletions). DeDrift (Baranchuk et al., 2023) handles data or query distribution shifts and identifies small and large partitions to re-cluster together, but it keeps the total number of partitions fixed. LIRE (Xu et al., 2023) performs localized split and merge operations when partitions grow too large or become too small after insertions and deletions. Ada-IVF (Mohoney et al., 2023) monitors partition size changes from updates and the partition access distribution from the queries. It combines this information to prioritize the maintenance operations that will improve index efficiency. Similar to us, these approaches demonstrate the effectiveness of localized maintenance operations for IVF indexes and the advantages of adapting to query distribution. However, they cannot be applied directly for our usage, which is incremental index construction as the number of queries increases. These works all trigger maintenance operations directly or implicitly from incoming updates (insertions/deletions). Thus they will never trigger in our static setting, whereas we do since we follow the database cracking paradigm, that states “index maintenance should be a byproduct of query processing, not of updates” (Idreos et al., 2007), and CrackIVF operations are triggered solely from the queries.
7. Conclusion and Future Work
We introduced CrackIVF, a cracking-based IVF index for Approximate Nearest Neighbor Search (ANNS), aimed to be used with RAG systems over embedding data lakes. By dynamically adapting to increasing query workloads, it eliminates the need for costly up-front indexing, allowing immediate query processing. Our experimental results show that CrackIVF can process over 1 million queries before conventional methods even finish building an index, achieving 10-1000x faster initialization times. This makes it particularly effective for cold data, infrequently accessed datasets, and rapid access to unseen data. As future work, we want to adapt CrackIVF to handle query and data distribution shifts and identify reliable methods to dynamically set parameters across diverse datasets.
8. Acknowledgments
We would like to thank AMD Research for the generous donation of the equipment in the ETHZ-AMD Heterogeneous Accelerated Compute Cluster (HACC) (https://systems.ethz.ch/research/data-processing-on-modern-hardware/hacc.html) that was used to run the experiments reported in this paper.

### v2

@@
 Retrieval Augmented Generation (RAG) uses vector databases to expand the expertise of an LLM model without having to retrain it.
-This idea can be applied over data lakes, leading to the notion of embeddings data lakes, i.e., a pool of vector databases ready to be used by RAGs.
+The idea can be applied over data lakes, leading to the notion of embedding data lakes, i.e., a pool of vector databases ready to be used by RAGs.
 The key component in these systems is the indexes enabling Approximated Nearest Neighbor Search (ANNS).
-However, in data lakes, one cannot realistically expect to build indexes for every possible dataset.
-In this paper, we propose an adaptive, partition-based index, CrackIVF, that performs much better than up-front index building.
-CrackIVF starts answering queries by near brute force search and only expands as it sees enough queries.
+However, in data lakes, one cannot realistically expect to build indexes for every dataset.
+Thus, we propose an adaptive, partition-based index, CrackIVF, that performs much better than up-front index building.
+CrackIVF starts answering as a small index, and only expands to improve performance as it sees enough queries.
 It does so by progressively adapting the index to the query workload.
 That way, queries can be answered right away without having to build a full index first.
-After seeing enough queries, CrackIVF will produce an index comparable to the best of those built using conventional techniques.
-As the experimental evaluation shows, CrackIVF can often answer more than 1 million queries before other approaches have even built the index and can start answering queries immediately, achieving 10-1000x faster initialization times.
-This makes it ideal when working with cold data or infrequently used data or as a way to bootstrap access to unseen datasets.
+After seeing enough queries, CrackIVF will produce an index comparable to those built with conventional techniques.
+CrackIVF can often answer more than 1 million queries before other approaches have even built the index, achieving 10-1000x faster initialization times.
+This makes it ideal for cold or infrequently used data and as a way to bootstrap access to unseen datasets.
 PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+PVLDB, 18(11): 3951 - 3964, 2025.
+doi:10.14778/3749646.3749666 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
 Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
@@
 Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
+Proceedings of the VLDB Endowment, Vol. 18, No. 11 ISSN 2150-8097.
+doi:10.14778/3749646.3749666
 PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
+The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/mageirakos/crack-ivf-vldb.
 1. Introduction
-External data sources can complement Large language models (LLMs) (Brown et al., 2020) through the use of retrieval-augmented generation (RAG) (Lewis et al., 2020).
-In RAG, external data is represented as vector embeddings (i.e., learned representations of data that result in a multidimensional vector).
-The embeddings are indexed by a variety of means and searched through approximate nearest neighbor (ANN) search so that LLMs can answer queries on information they were not explicitly trained on.
+Large language models (LLMs) (Brown et al., 2020) can be complemented with retrieval-augmented generation (RAG) (Lewis et al., 2020).
+In RAG, external data is represented as vector embeddings (i.e., learned representations of data that results in a multidimensional vector).
+It is indexed and accessed through approximate nearest neighbor (ANN) search so that LLMs can answer queries on information they were not trained on.
 The effectiveness of RAG is tied to that of the ANN index structures, making them a critical component.
-LLMs have been integrated with search engines (OpenAI, 2024) (e.g., Google, Bing, and Baidu).
-However, the vast repository of private, unstructured business data remains largely untapped.
-It has been estimated that unstructured data constitutes up to 80-90% of the total data volume (Merrill Lynch, 1998; MIT Sloan School of Management, 2021) and is often stored in open data formats on data lakes (Miller, 2018).
-However, as much as 70% of this data (Data Science Association, 2016) remains unused, classified as “dark” data (Heidorn, 2008; Zhang et al., 2016).
-This data can be made discoverable (Chapman et al., 2020), but even then, the actual data often remains unindexed.
-For instance, as of March 2020, Google Research’s Dataset Search (Brickley et al., 2019) had indexed around 28 million structured and unstructured datasets (Benjelloun et al., 2020) based on their metadata.
-However, embedding-based vector search and RAG techniques make the promise to go a step further by enabling direct retrieval and question answering on the underlying data itself, inside of each dataset.
-This requires embedding the data, and the creation of approximate nearest neighbor (ANN) indexes for every single dataset, exposing the data to RAG systems.
-In this paper, we name such an approach embedding data lakes (EDL), where unstructured data is stored alongside its vector representations.
-Efforts in this direction are already ongoing in both industry and academia.
-Databricks has deployed vector search capabilities (Databricks, 2024) over their Lakehouse (Armbrust et al., 2021), and new custom storage formats like Lance (LanceDB Developers, 2024) are being used to store embeddings on data lakes.
-Researchers are also directly exploring RAG-based techniques to query multi-modal data lakes (Tang et al., [n.d.]; Chen et al., 2023), as well as answering queries directly over unstructured data (Liu et al., 2025; Anderson et al., 2024).
-There is also work to directly define LLM-based data discovery systems over data lakes (An et al., 2025).
-The issue we address in this paper is crucial to achieve the promise of embedding lakes (EDL): index selection at scale.
-EDL, in principle, would require building millions of ANN indexes across diverse datasets, modalities, workloads, and embeddings, making static index selection and pre-computation infeasible.
-In data lakes, some datasets may receive millions of queries, while others are rarely accessed.
+Figure 1. Total time to answer the queries submitted for different indexing strategies vs the number of queries submitted
+In practice, the vast majority of private, unstructured data remains untapped.
+It is estimated to be up to 80-90% of all data (Merrill Lynch, 1998; Harbert, 2021), and is often stored in data lakes on open data formats (Miller, 2018).
+As much as 70% of it remains unused (Gualtieri, 2016), classified as “dark” data (Heidorn, 2008; Zhang et al., 2016).
+While the datasets can be made discoverable (Chapman et al., 2020), the actual underlying data in them often remains unindexed.
+As of March 2020, Google Research’s Dataset Search (Brickley et al., 2019) had indexed around 28 million structured and unstructured datasets (Benjelloun et al., 2020) based on their metadata.
+Embedding-based vector search and RAG techniques could however help us go a step further by enabling direct retrieval and question answering on the underlying data itself.
+This requires embedding the data, as well as the creation of approximate nearest neighbor indexes so as to expose the data to RAG systems.
+We refer to such an approach as embedding data lakes (EDL), where unstructured data is stored alongside its vector representations and is queried by a pool of vector databases.
+Efforts in this direction are already ongoing.
+Databricks uses vector search (Databricks, 2024) over their Lakehouse (Armbrust et al., 2021), and new custom storage formats (LanceDB Developers, 2024) are being used to store embeddings on data lakes.
+Researchers are also exploring RAG-based techniques to query multi-modal data lakes (Tang et al., 2024; Chen et al., 2023), as well as answering queries over unstructured data (Liu et al., 2025; Anderson et al., 2024).
+The issue we address in this paper is index selection at scale.
+Embedding data lakes, in principle, would require building millions of indexes across diverse datasets, modalities, workloads, and embedding models, which is neither feasible nor cost efficient.
+Some datasets may receive the majority of queries while others are rarely accessed.
 When selecting an index for embedding lakes, there is a trade-off.
-On the one hand, building an index requires significant upfront costs that might not pay off if the data is rarely used.
+On the one hand, building an index requires significant upfront costs that might not pay off if the data is rarely queried.
 On the other hand, skipping the indexing process and relying on a brute-force search is not scalable.
-Hence, data must be indexed using Approximate Nearest Neighbor (ANN) indexes but this raises two questions: when to index the data and how to index it.
-Figure 1. Total time to answer the queries submitted for different indexing strategies vs the number of queries submitted
-An optimal ANN index depends on workload patterns and future usage, which are often unknown.
+Hence, data must be indexed, which raises two questions: when and how to index the data.
+An optimal ANN index depends on the workload patterns and future usage, which are often unknown.
 As shown in Figure 1, the index that minimizes total cumulative time spent on building and searching varies with the query volume.
 Larger indexes take longer to build and require many queries to justify their upfront cost (e.g., the large IVFlat configurations (5000, 16000 clusters) in the figure).
-Brute-force search (as shown in the figure), and smaller indexes (IVFlat 1000 clusters), minimize time-to-query for fast discovery but struggle at scale due to their higher query response times.
+Brute-force search, and smaller indexes (IVFlat 1000 clusters), minimize time-to-query for fast discovery but struggle at scale due to their higher query response times.
 To tackle index selection at scale, we adopt a well-established strategy: carefully avoiding the need to build the full index upfront.
-In this, we draw inspiration from deferred data structuring (Karp et al., 1988) and database cracking (Idreos et al., 2007; Halim et al., 2012; Schuhknecht et al., 2016), which have been successfully applied to single-dimensional database indexing.
+We draw inspiration from deferred data structuring (Karp et al., 1988) and database cracking (Idreos et al., 2007; Halim et al., 2012; Schuhknecht et al., 2016).
 Similar approaches have been extended to multidimensional data (Holanda et al., 2018; Zardbani et al., 2023; Pavlovic et al., 2018; Nerone et al., 2021).
-However, prior work does not scale to the embedding dimensionalities nor supports the k-Nearest Neighbour (k-NN) queries typical in RAG systems.
-Recently, the AV-Tree was the first adaptive index for high-dimensional k-Nearest Neighbor (k-NN) search (Lampropoulos et al., 2023).
+However, prior work does not scale to the embedding dimensionalities nor supports the k-Nearest Neighbour (k-NN) queries typical in RAG deployments.
+Recently, the AV-Tree was the first adaptive index for high-dimensional k-NN search (Lampropoulos et al., 2023).
 AV-Tree targets short-lived data, with workloads of up to one thousand queries, and is a tree-based Exact Nearest Neighbor (ENN) adaptive index.
-This makes it impractical for RAG-scale workloads, which rely on Approximate Nearest Neighbor (ANN) index structures to trade off a bit of accuracy for large performance gains.
-Nonetheless, the idea of dynamically building an index tailored to the workload remains compelling.
-In this paper, we demonstrate how to achieve this efficiently in the context of RAG and employing the same data structures (Sivic and Zisserman, 2003) and systems (Douze et al., 2024) used for ANN search workloads (Aumüller et al., 2020; Simhadri et al., 2022) rather than analytics over cold data.
+This makes it impractical for RAG-scale workloads, which rely on Approximate Nearest Neighbor index structures to trade off a bit of accuracy for lower query response time and higher scalability.
+Nonetheless, the idea of dynamically building an index tailored to the workload is compelling.
+In this paper, we demonstrate how to do this efficiently in the context of RAGs by employing the same data structures (Sivic and Zisserman, 2003) and systems (Douze et al., 2024) used for ANN search workloads (Aumüller et al., 2020; Simhadri et al., 2022).
 We introduce CrackIVF, an incrementally built, partition-based ANN index that dynamically improves itself to adapt to increasing workload demands, minimizing both time-to-query and cumulative cost associated with suboptimal static index selection.
 CrackIVF evolves as a side effect of query execution.
-Each query is a candidate to add a new partition, or crack, while its search region is a candidate for local refinement, allowing the index to grow and improve as more queries arrive.
-Efficiency is maintained through two controls: one deciding where to apply cracking and refinement in the search space, and another controlling when, to balance indexing and search.
-In the paper we focus on a static setting (i.e., no data or query distribution shifts) and evaluate CrackIVF on multiple standard open source datasets for ANN search (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016).
+Each query is a candidate around which we can add a new partition, or crack.
+Likewise, its local region visited during search is a candidate for local refinement, allowing the index to grow and improve as more queries arrive.
+Efficiency is maintained through two control mechanisms: one deciding where to apply cracking and refinement in the search space, and another controlling when, balancing indexing and search times.
+In this work we focus on the static setting (i.e., no data or query distribution shifts) and evaluate CrackIVF on multiple standard open source datasets for ANN search (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016).
 For partition-based indices, index selection means deciding upfront how many partitions the index will use, which has an impact on the ANN search performance.
+More partitions increase the upfront construction cost but improve query performance.
 This decision has to be made without knowing the future query workload.
-CrackIVF waits to build the index until it sees enough queries and decides on the clusters based on the query distribution over the search space.
+On the other hand, CrackIVF waits to build the index as it sees enough queries and decides on the clusters based on the query distribution over the search space.
 As more queries arrive, the clusters are redefined, and their number is increased to adapt to the increasing query workload.
-Asymptotically, CrackIVF builds an index that converges to similar to those pre-built, but it has been able to answer queries along the way.
+CrackIVF is an incrementally built index that asymptotically matches or surpasses the performance of pre-built indexes, while it has been able to answer queries along the way.
 Across benchmarks, CrackIVF consistently outperforms pre-built partition-based indexes.
 It converges to near-optimal query response time, yet achieves several orders of magnitude lower startup time than other indexes.
-It efficiently scales to large workloads and, in some cases, it can process 1 million queries before the baseline indexes have finished building.
+It efficiently scales to large workloads and, in some cases, can process 1 million queries before the baseline indexes have finished building.
 It is the only index that consistently remains near the Pareto frontier of minimum cumulative time across the number of queries posed to the system.
 2. Background
-In this section, we briefly cover the basics of approximated nearest neighbor search and IVF indexes.
 Nearest Neighbor Search In the k-nearest neighbor (k-NN) search problem, we are given a set of points $P\in\mathbb{R}^{|P|\times d}$ and a query $q\in\mathbb{R}^{d}$.
-The aim is to find the $k$ points in $P$ that are closest to $q$ under some notion of distance or similarity (e.g., minimize the Euclidean distance (L2) or maximize the inner product (IP) to measure closeness).
-A straightforward Exact Nearest Neighbor (ENN) approach is a linear scan of every point, a.k.a brute-force search, incurring in $\mathcal{O}(|P|\cdot d)$ complexity.
-There exist more sophisticated exact indexing structures (e.g., KD-trees (Bentley, 1975), R-trees (Guttman, 1984), or VP-trees (Yianilos, 1993)).
-However, they often degrade to near-linear-scan performance in higher dimensions, due to the curse of dimensionality (Weber et al., 1998), making them impractical when $|P|$ or $d$ are large.
-With embeddings of hundreds to thousands of dimensions  (OpenAI, 2023) and real-world datasets at the scale of billions of data points (Simhadri et al., 2022), RAG systems turn to Approximate Nearest Neighbor (ANN) search techniques.
+The aim is to find the $k$ points in $P$ that are closest to $q$ under some notion of distance or similarity (e.g., minimize the Euclidean distance (L2) or maximize the inner product (IP)).
+A straightforward Exact Nearest Neighbor approach is a linear scan of every point, a.k.a brute-force search, incurring $O(|P|\cdot d)$ complexity.
+There exist more sophisticated exact indexing structures (e.g., KD-trees (Bentley, 1975), R-trees (Guttman, 1984)).
+However, due to the curse of dimensionality (Weber et al., 1998), performance degrades to near-linear-scan in higher dimensions, making them impractical when $|P|$ or $d$ are large.
+With embeddings of hundreds to thousands of dimensions  (OpenAI, 2025) and real-world datasets at the scale of billions of data points (Simhadri et al., 2022), RAG systems turn to Approximate Nearest Neighbor search techniques.
 They allow a small retrieval error $\varepsilon$ in exchange for better runtime or memory usage.
-This approximation quality can also be evaluated by the fraction of exact top-$k$ vectors recovered.
+The approximation quality can be evaluated by the fraction of exact top-$k$ vectors recovered.
 (a) Response Time with increasing partitions and training K-Means.
@@
 Figure 2.
-Observations behind the design: (a) separating number of partitions and refinement of centroids; (b) access to the index is generally skewed towards certain regions; (c) regions queried often can be clustered and refined at a much lower granularity
+Observations behind the design: (a) separating the number and refinement of partitions; (b) access to the index is generally skewed towards certain regions; (c) regions queried often can be clustered and refined at a much lower granularity
 Partition based IVF methods One of the most common ANN search structures is partition-based IVF indexes (Sivic and Zisserman, 2003; Chen et al., 2021; Sun et al., 2024).
@@
 Since the remaining partitions are skipped, some nearest neighbors may be missed, which is why the result is an approximate solution.
-A partition-based index is built by first selecting $n_{\text{list}}$ partition representative vectors, typically using $k$-means on a training subset $P_{\text{train}}\subset P$.
-Each data point is then assigned to the nearest representative, forming disjoint partitions.
+A partition-based index is built by first selecting $n_{\text{list}}$ representative vectors for each partition, typically using $k$-means on a training subset $P_{\text{train}}\subset P$.
+Each data point is then assigned to the nearest representative.
 These assignments are typically stored in an inverted file (IVF) structure (Sivic and Zisserman, 2003).
-A widely adopted implementation of IVF indexes is the FAISS library from Meta (Douze et al., 2024), which we use as a baseline.
+A widely adopted implementation of IVF indexes is the FAISS library from Meta (Douze et al., 2024), which we built CrackIVF on and use as a baseline.
 See Section 6 for other ANN index types.
 3. Solution Overview
-We focus on Partition-based IVF indexes because they offer several advantages for deployment in embedding data lakes.
-They offer the best trade-off in index build time, query response time, and index size (Research, 2023).
-They have the lowest index build times relative to query throughput (QPS).
-All of the above are crucial for embedding data lakes, where potentially millions of indexes of various sizes and levels of utilization will have to be constructed and stored.
+[…]
 3.1. Cracking in high dimensions
-A well known approach to avoid upfront cost for indexing or organizing data is cracking.
-In database cracking (Idreos et al., 2007), there is a final state where a single column is sorted.
-Over time, fewer points are moved, and over shorter distances, reducing the overhead of the cracking.
+An approach to avoid the upfront cost for indexing is cracking.
+In the original database cracking (Idreos et al., 2007), the final state is where a single column is sorted.
+Over time, fewer points are physically reorganized reducing the overhead.
+In high-dimensional spaces, the having a final sorted state no longer holds, and space filling curves, like Z-order (Samet, 1984), do not scale past a few dimensions.
 Additionally, in a naive implementation, the physical reorganization happens with every incoming query.
-In high-dimensional spaces, the assumption of a fully sorted state no longer holds, and space filling curves, like Z-order (Samet, 1984), do not scale past a few dimensions.
-Additionally, enforcing reorganization after every query becomes increasingly difficult, as it is a more expensive operation.
-Moving points between inverted lists incurs random access overhead, with no eventual locality benefits.
+In high-dimensions, this becomes increasingly difficult.
 High-dimensional vectors require moving multiple bytes per dimension, increasing data movement costs.
-Furthermore, inverted lists, are often implemented as memory-aligned contiguous arrays to allow for efficient SIMD-based kernels, and they may require linear-time compaction even for a single vector to move.
-These factors contribute to the significant latency overhead, making physical reorganization after every query difficult.
-Based on this, the challenge to address is as follows A cracking-based IVF index must efficiently manage the significantly higher overhead of physical reorganization in high dimensions, while minimizing time-to-query, continuously improving the index, and ensuring minimum cumulative time across all scales.
+In IVF, moving points between inverted lists incurs random access overhead, with no eventual locality benefits.
+Furthermore, inverted lists are often implemented as memory-aligned contiguous arrays to allow for efficient SIMD-based operations, and they may require linear-time compaction even for a single vector to move.
+These factors contribute to the latency overhead, making physical reorganization after every query difficult.
+The challenge is as follows: A cracking-based IVF index must efficiently manage the significantly higher overhead of physical reorganization in high dimensions, while minimizing time-to-query, continuously improving the index, and ensuring minimum cumulative time across all scales.
 3.2. Outline of the approach
-As Figure 2(a) illustrates, index construction can be divided into two operations that independently improve index performance: (i) increasing the number of partitions by randomly selecting partition representatives and (ii) refining these representatives into well-placed centroids using k-means training, which helps to distribute points more evenly across partitions and improve the overall representativeness.
-Although both improve the index independently, the best results consistently come after reaching a maximum index size, even with randomly initialized representatives, followed by k-means training to optimize their placement.
-Thus, the optimal approach is to first increase the number of partitions up to a limit and then refine their locations using k-means.
-Observation 1 behind the approach: Index construction can be decoupled into two distinct build operations, which can improve the index independently and be separated in time. .
-Existing partition-based IVF vector search indexes apply build operations evenly across the vector space, with centroids sampled uniformly and K-means refinements performed simultaneously throughout.
-While simple and intuitive, this approach assumes equal importance across the space, and requires that both operations are applied at the same time on the entire space leading to larger than necessary index build times.
+Index construction can be divided into two operations that independently improve index performance, i) increasing the number of partitions and ii) refining partitions by using K-means, which helps to distribute points more evenly (Figure 2(a)).
+The best results consistently come after reaching a maximum index size, even with randomly chosen representatives, followed by K-means training to optimize their placement.
+Thus, the optimal approach is to first increase the number of partitions up to a limit and then also refine their locations using K-means.
+Observation 1: Index construction can be decoupled into two distinct build operations, which can improve the index independently and be separated in time.
+Existing partition-based IVF vector search indexes apply build operations evenly across the vector space, with partitions sampled uniformly and K-means refinements performed globally.
+While simple and intuitive, this approach assumes equal importance across the space, leading to larger than necessary index build times.
 However, vector search workloads can exhibit highly skewed access patterns, where certain regions are significantly more important than others.
-For instance, in an industrial workload (Mohoney et al., 2023, 2024), up to 85% of centroids are never accessed during search, which would make partitioning or training efforts in those regions unnecessary.
+For instance, in an industrial workload (Mohoney et al., 2023, 2024), up to 85% of partitions are never accessed during search, which would make partitioning or training efforts in those regions unnecessary.
 As we show in Figure 2(b), skewed query distributions also appear in well known open-source vector search datasets (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016).
 Even on open source datasets showcasing moderate amounts of skew, where $80\%$ of the total queries access $60\%$ of the partitions, the disparity between the most frequently and least frequently accessed partitions is upwards of 50 times.
-There is a fundamental inefficiency in current uniform indexing approaches: a large portion of the vector space may rarely accessed, yet the indexing effort is evenly distributed across the space, and always paid upfront.
-Observation 2 behind the approach: Just as index construction can be separated in time, it can also be separated in space.
-The index build operations do not have to be applied uniformly across the vector space; instead, they can be focused on regions that are accessed more frequently.
-Since certain areas of the vector space are more critical than others, adapting index construction to follow the query distribution is a natural strategy to optimize efficiency and accelerate indexing.
-Lastly, in partition-based IVF indexes, a query first identifies the nearest partitions and then scans all points within those partitions to retrieve the $k$ nearest neighbors.
-This involves the same distance computations needed to update the index, which naturally leads to a cracking-based approach, where index refinement occurs as a side effect of query execution.
-Observation 3 behind the approach: We can further amortize the cost of specific build operations, such as assigning vectors to partitions and identifying the local regions to crack and refine, Figure 2(c).
+There is a fundamental inefficiency in current uniform indexing approaches.
+A large portion of the vector space may rarely accessed, yet the indexing effort is evenly distributed across the space, and always paid upfront.
+Observation 2: Just as index construction can be separated in time, it can also be separated in space.
+Index build operations do not have to be applied uniformly across the vector space.
+Instead, they can be local to regions that are accessed.
+Adaptive index construction that follows the query distribution is one strategy to optimize for efficiency.
+In partition-based IVF indexes, operations performed during search and those performed during index construction are very similar.
+During search the distances of the query to centroids and nearby points are computed, and during index construction the distances of points to centroids are computed.
+This means that by treating the query as a potential new partition representative “centroid”, we can both identify good locations to add new cracks, and reuse the intermediate computations of search to perform incremental index construction.
+This naturally leads to a cracking-based approach for IVF, where index construction can occur as a side effect of query execution.
+Observation 3: We can further amortize the cost of specific build operations, such as assigning vectors to partitions and identifying the local regions to crack and refine, Figure 2(c).
 We can reuse computations from search operations to perform index construction as a side effect of query execution.
 3.3. Outline of the solution
-Based on these insights, we propose a partition-based crackable IVF index for ANN search.
-It is designed to remain near the optimal minimum cumulative time across the number of queries submitted (Figure 1), improving as the workload increases, and addressing the core issue of indexing parts of the search space that are not used.
-CrackIVF employs two localized and independent build operations: CRACK and REFINE.
-The former introduces new partitions, while the latter applies a localized K-means to improve them.
-These operations run independently across both time and vector space.
-We directly follow the query distribution, where queries themselves act as candidates for CRACK and REFINE, determining where new partitions should be introduced, so they are not randomly sampled, and which local regions need refinement.
-Intermediate results from search operations, such as the distances computed to nearby points, can be re-used to amortize the cost of assignments.
-Using the queries to create partitions, rather than selecting random ones, allows distances computed during search to reused, steal points from the local region, that are closer to the query than their current assignment.
-Additionally, using the queries as candidates for CRACK and REFINE enables CrackIVF to exactly follow the query distribution in terms of when and where to execute build operations, as partitions that receive a larger number of queries will proportionally be candidates for cracking and refinement by the same amount.
+We propose a partition-based crackable IVF index for ANN search.
+It is designed to remain near the minimum cumulative time across the number of queries submitted, be constructed as the query load increases, and avoid indexing unused parts of the vector space.
+At initialization, CrackIVF starts as a small IVF index, with coarse partitioning of the entire dataset.
+As queries come in, more partitions are added in the accessed regions related to the queries (Figure 2(c)).
+The index employs two localized and independent build operations: CRACK and REFINE.
+The former introduces new partitions, while the latter applies a localized K-means to refine them.
+These operations run independently across both time and region of the vector space.
+Queries and their local visited regions act as CRACK and REFINE candidates, determining where new partitions should be introduced.
+Using the queries to create partitions, rather than selecting random ones, allows the distances computed during search to be used for stealing points from the local region, that are closer to the query than their current assignment.
+Thus, intermediate results from search operations are reused to amortize the cost of construction.
+Additionally, using the queries as candidates for CRACK and REFINE enables CrackIVF to follow the query distribution in terms of when and where to execute build operations.
+Partitions that receive a larger number of queries will be proportionally more likely to be selected for cracking and refinement.
 Finally, to mitigate the high overhead of storage reorganization in high-dimensional IVF indexes, we do not perform build operations after every query.
-Instead, CRACK operations are buffered and committed when beneficial, while REFINE is executed eagerly but infrequently to improve local regions that queries are visiting when necessary.
-To automate these decisions, we introduce two control mechanisms.
-The first is a set of heuristic rules, that determine where in the vector space cracks should be buffered or refinements executed.
+Instead, CRACK operations are buffered and committed only when it is beneficial, while REFINE operations are executed eagerly, but infrequently, to improve local regions that queries are visiting.
+We introduce two control mechanisms to make these decisions.
+The first is a set of heuristic rules that determine the decision boundaries for where in the vector space cracks should be buffered or refinements executed.
 The second is a cost-based approach that estimates the latency overhead of a build operation and constrains it relative to the cumulative runtime of the index since initialization.
-As more queries arrive and the index is used more frequently, the budget for build operations increases proportionally, allowing the index to grow and refine in sync with the query workload.
+As more queries arrive and the index is used more frequently, the budget for build operations increases proportionally, allowing the index to improve as the query workload demands increase.
+Table 1. Notation used in algorithms and cost model.
 Symbol Description
 $|X|$ Cardinality, e.g., $|X|=N$ if $X\in\mathbb{R}^{N\times d}$
-[…]
+$metric(a,b)$ Similarity metric, where in our work
+$metric\in\{\text{Eucl. Distance (L2)},\text{Inner Product (IP)}\}$
 $P\in\mathbb{R}^{|P|\times d}$ Data points, $d$-dimensional
 $C\in\mathbb{R}^{|C|\times d}$ Committed cracks
-$C_{\text{local}}\in\mathbb{Z}^{nprobe}$ Local region centroid IDs
+$C_{\text{local}}\in\mathbb{Z}^{nprobe}$ Local region crack IDs
+$P_{\text{local}}$ Local region points
 $P_{\text{train}}$ K-means training points, where $|P_{\text{train}}|=\min(|P_{\text{local}}|,|C_{\text{local}}|\times\max_{\text{points}})$
 $n_{\text{iter}}$ K-means iterations
-$\max_{\text{points}}$ K-means max training sample per centroid
-$nlist$ Total inverted lists storing $P$, where $nlist=|C|$
+$\max_{\text{points}}$ K-means max training sample per crack
+nlistnlist Total inverted lists storing $P$, where $nlist=|C|$
 $Q\in\mathbb{R}^{b_{s}\times d}$ Query set per search
 $D\in\mathbb{R}^{b_{s}\times|r|},\,I\in\mathbb{Z}^{b_{s}\times|r|}$ kNN distances and point indices from SEARCH, where $|r|\in[k,|P_{\text{visited}}|]$
-$bs$ Total queries batched in single SEARCH
-$nprobe$ Number of partitions scanned per SEARCH
+bsbs Total queries batched in single SEARCH
+nprobenprobe Number of partitions scanned per SEARCH
 $k$ Number of nearest neighbors to return
-$A_{P}^{\ast}\in\mathbb{Z}^{|P|}$ Point assignments to centroid IDs
-$D_{P}^{\ast}\in\mathbb{R}^{|P|}$ Point distances to centroids
-$H_{C}^{\ast}\in\mathbb{N}^{|C|}$ Histogram of points per centroid
-[…]
-Table 1. Notation used in algorithms and cost model.
+$\alpha$ Ratio of total build time to total time
+$A_{P}^{\ast}\in\mathbb{Z}^{|P|}$ Point assignments to crack IDs
+$D_{P}^{\ast}\in\mathbb{R}^{|P|}$ Point distances to crack representative
+$H_{C}^{\ast}\in\mathbb{N}^{|C|}$ Histogram of points per crack
+Note: $\ast\in\{\text{dyn},\text{true}\}$, where dyn refers to the dynamic state (includes buffered cracks) and true refers to the true index state.
 4. Crack-IVF
-CrackIVF, is a partition-based crackable IVF index implemented in FAISS (Douze et al., 2024).
-The core logic of the algorithm and operations are covered in Section 4.1 and Algorithm 1.
-Control mechanisms, including heuristic rules and cost models, are described in Section 4.2 and Section 4.3.
-They aid in making accurate decisions about where and when to perform the build operations, CRACK and REFINE, which are shown in Algorithm 2 and Algorithm 3.
-The notation used in this section is summarized in Table 1.
+CrackIVF, is implemented in FAISS (Douze et al., 2024).
+We describe the core logic of the algorithm and operations in Section 4.1 and Algorithm 1; the control mechanisms, including heuristic rules and cost models, in Sections 4.2 and 4.3.
+These aid in making accurate decisions about where and when to perform the build operations, CRACK and REFINE (Algorithms 2 and 3).
+The notation used is in Table 1.
 4.1. Algorithm and Operations
-CrackIVF performs index construction and storage reorganization operations as a side-effect of query execution to incrementally improve index performance.
-It has three core operations, $SEARCH$ finds the k-NN results to return to the user, and $CRACK$, $REFINE$ are the build operations that perform physical reorganization of the IVF structure as well as change the index state.
-SEARCH: The search procedure is common among partition based IVF indexes, where the local region to the query is identified by the nearest partitions, whose point assignments are subsequently scanned to find the final $k$ nearest neighbors of the query.
-In our case, the result are point ids, $I$, and distances $D$ to the query, of the visited points in the local region sorted by similarity, together with the ids of the cracks that the query visited, $C_{visited}$, as seen in Line 7 of Algorithm 1.
+CrackIVF performs index construction and storage reorganization operations as a side-effect of query execution to incrementally improve the index.
+It has three core operations, $SEARCH$ finds the k-NN results to return to the user, while $CRACK$ and $REFINE$ operations reorganize the IVF structure and change the index state.
+SEARCH: During the search procedure of IVF, the local region to the query is identified by finding the nprobenprobe nearest partitions to the query.
+Subsequently, the point assignments to these local partitions are scanned to find the $k$ nearest neighbors.
+The result includes the nearest point ids $I$, their distances $D$ to the query, and the $C_{visited}$ ids of the local region cracks that the query visited.
 Algorithm 1 SearchAndCrack
@@
 7:   $(D_{\text{local}},I_{\text{local}},C_{\text{local}})\leftarrow(D[q],I[q],C_{visited}[q])$
-8:     $I_{\text{steal}}\leftarrow\{p\in I_{\text{local}}\mid D_{\text{local}}(p)\mathrel{\text{op}}D_{P}^{\text{dyn}}(p)\}$op = $<$ for L2, $>$ for IP}
-9:     $good\_crack\_candidate$ = CrackHeuristic($I_{steal},\text{State}_{\text{dyn}}$)
-10:     if $good\_crack\_candidate$ then
+8:   $I_{\text{steal}}\leftarrow\{p\in I_{\text{local}}\mid D_{\text{local}}(p)\mathrel{\text{op}}D_{P}^{\text{dyn}}(p)\}$ {op = << for L2, >> for IP}
+9:   good_crack_candidategood\_crack\_candidate = CrackHeuristic($I_{steal},\text{State}_{\text{dyn}}$)
+10:   if good_crack_candidategood\_crack\_candidate then
 11:    $C_{\text{buffered}}\leftarrow C_{\text{buffered}}\cup\{q\}$
@@
 16:  \\ Step 3: CRACK Decision (When to Commit, see §4.3)
-[…]
+17:  $\hat{T}_{\text{CRACK}}\leftarrow\textsc{EstimateCrackCost}()$
 18:  $can\_afford\_crack\leftarrow(T_{\text{build}}+\hat{T}_{\text{CRACK}}\leq\alpha\cdot(T_{\text{build}}+T_{\text{search}}+\hat{T}_{\text{CRACK}}))$
 19:  $enough\_buffered\leftarrow(|C_{\text{buffered}}|>C_{\min})$
-20:  if $can\_afford\_crack$ and $enough\_buffered$ then
+20:  if can_afford_crackcan\_afford\_crack and enough_bufferedenough\_buffered then
 21:   \\ UNDO if REFINE invalidated previously good cracks
@@
 25:   $T_{\text{start\_crack}}\leftarrow\textsc{CurrentTime}()$
-26:     CRACK$(C_{\text{buffered}},I_{\text{buffered}})$Algorithm 2}
+[…]
 27:   $T_{\text{build}}\leftarrow T_{\text{build}}+(\textsc{CurrentTime}()-T_{\text{start\_crack}})$
@@
 31:   for each query $q\in\textit{Q}$ do
-[…]
+32:    $\hat{T}_{\text{REFINE}}\leftarrow\textsc{EstimateRefineCost}()$
 33:    $can\_afford\_refine\leftarrow(T_{\text{build}}+\hat{T}_{\text{REFINE}}\leq\alpha\cdot(T_{\text{build}}+T_{\text{search}}+\hat{T}_{\text{REFINE}}))$
-34:        if $can\_afford\_refine$ then
+34:    if can_afford_refinecan\_afford\_refine then
 35:     $good\_refine\_candidate\leftarrow\textsc{RefineHeuristic}(C_{\text{local}},\text{State}_{\text{true}})$
-36:           if $good\_refine\_candidate$ then
+36:     if good_refine_candidategood\_refine\_candidate then
 37:      $T_{\text{start\_refine}}\leftarrow\textsc{CurrentTime}()$ {Start timing REFINE}
-38:              REFINE()($\dots$)( … )
+38:      REFINE()($\dots$)
 39:      $T_{\text{build}}\leftarrow T_{\text{build}}+(\textsc{CurrentTime}()-T_{\text{start\_refine}})$
@@
 CRACK: Physical reorganization of points between inverted lists, and the mutation of the index does not happen after each query.
-Each incoming query represents a crack candidate, which is evaluated by heuristic rules to classify it as good or bad (Section 4.2) (Line 9 Algorithm 1).
+Each incoming query represents a crack candidate, which is evaluated by rules that define the decision boundaries which classify it as good or bad (Section 4.2) (Line 9 Algorithm 1).
 Only good cracks are further buffered and potentially committed at a later time.
 Each crack is essentially a new partition, where the points assigned to the partition have been “stolen” based on their proximity to the query.
-We reuse the results of the SEARCH operations to compare if the distance of a point to the query is closer than the distance of the point to its current assignment, which could be another buffered crack.
-If it is, then it’s beneficial to reassign the point to the new crack, and this assignment is also buffered, to be committed when $CRACK$ is executed, as explained in Section 4.3.
-$CRACK$ is the procedure performing the physical reorganization, syncing the buffered state of the index with the currently working state.
-The core individual kernels of the procedure are sequentially executed (Algorithm 2).
+We reuse the results of the SEARCH operations to compare if the distance of a point to the query is closer than the distance of the point to its current assignment.
+If it is, then it’s beneficial to reassign the point to the new crack, and this assignment is buffered, to be committed when $CRACK$ is executed, as explained in Section 4.3.
+$CRACK$ is the procedure that performs the physical reorganization, synchronizing the buffered state of the index with the currently working state.
+The core individual kernels of the procedure are shown in Algorithm 2.
 It adds a new inverted list for each buffered crack, moves stolen points to their new assignments, and updates the partition representatives to be the centroid of the points based on their new assignments.
-This updates the location of the crack representatives, whose point assignments were stolen, in order for their new location to match the underlying assignment distribution.
-This procedure is only executed when enough budget for build operations has been accumulated through the usage of the index (Line 18 Algorithm 1).
-This control mechanism ensures that CrackIVF does not spend a disproportionate amount of time on CRACK operations.
-The parameter $\alpha$, the ratio of total build operations time to total search time, is what controls the available budget.
-Thus, since $CRACK$ is lazy, the heuristics governing where a crack is added and the cost-based control mechanism for when it is added are independent.
+This procedure is only executed when enough budget for build operations has been accumulated throughout the usage of the index (Line 18 Algorithm 1).
+This budget control mechanism ensures that CrackIVF does not spend a disproportionate amount of time on CRACK operations.
+The parameter $\alpha$, the ratio of total build operations time to the total time, is what controls the available budget.
+Since $CRACK$ is applied lazily, the heuristics governing where a crack is added and the cost-based control mechanism for when it is added are independent.
 Algorithm 2 CRACK
@@
 4:  Update Index State: Synchronize $State_{dyn}$ and $State_{true}$ to reflect the changes.
-REFINE: Refinement is the second operation to improve the index.
+REFINE: Refinement is the second operation that improves the index.
 It is performed eagerly, focusing on the local region visited by the query.
-The goal is to immediately correct suboptimal regions in the current index.
-Thus, the buffered states and buffered cracks within the local region are not considered.
-The decision to refine a region is based on heuristic rules that evaluate the imbalance of point assignments in local cracks.
-Additionally, executing a refinement operation depends on an estimate of the runtime cost, determined using a cost model specific to $REFINE$ (Line 33 Algorithm 1).
-Unlike cracking, the opportunity to refine a local region is transient.
-The heuristic rules must decide that the currently visited local region is a good candidate for refinement, and at the same time, there must be sufficient computational budget to execute the operation.
-If where and when to execute a refine do not simultaneously agree, the next opportunity for the same region will arise only when another query accesses it.
-$REFINE$, as described in Algorithm 3, executes a sequence of kernels implementing a local K-means variation and performs physical reorganization.
+A local region is defined by the nprobenprobe nearest partitions that the query visits during search.
+Algorithm 1 Line 7, is where the local region $C_{local}$ for a query is extracted from the SEARCH result $C_{visited}$ of a query batch.
+The goal of REFINE is to immediately correct suboptimal assignments in the regions of the current index.
+Thus, any buffered cracks and assignments within the local region are not considered, as they are not part of the current index state.
+The decision to refine a region is based on heuristics evaluating the imbalance of point assignments in local cracks.
+Executing a refinement operation depends on an estimate of the runtime cost, determined using a cost model specific to $REFINE$ (Line 33, Algorithm 1).
+Unlike cracking, the opportunity to refine a local region takes place when both the currently visited local region is a good candidate for refinement and there is sufficient computational budget to execute the operation.
+If the “where” and “when” heuristics for $REFINE$ do not simultaneously agree, the next opportunity is when another query accesses the same region.
+$REFINE$ is shown in Algorithm 3).
+It executes a sequence of kernels implementing a local K-means variation and performs physical reorganization.
 This process improves the placement of crack representatives within the local region and reassigns points to their nearest representative.
-Unlike $CRACK$, $REFINE$ incurs additional computational costs as it does not rely on precomputed buffered assignments.
-Instead, assignments are determined dynamically upon completion of the local K-means process.
+$REFINE$ has higher computational costs as it does not rely on precomputed buffered assignments, and assignments are determined dynamically upon completion of the local K-means process.
+$CRACK$ has a higher data-movement cost, since typically a larger region of the vector space is affected when cracks are added throughout.
 Algorithm 3 REFINE
@@
 1:  Get Local Region: Retrieve all points $P_{local}$ in local region.
-2:  Local K-Means: Refine region of $C_{\text{local}}$, producing $C^{\prime}_{\text{local}}$ centroids.
+2:  Local K-Means: Refine region of $C_{\text{local}}$, producing $C^{\prime}_{\text{local}}$.
 3:  Commit Reorg (w/o $\mathbf{A_{P}^{\text{dyn}}}$): Replacing $C_{\text{local}}$ with refined $C^{\prime}_{\text{local}}$.
@@
 4:  Update Index State: Synchronize $State_{dyn}$ and $State_{true}$ to reflect the changes.
-Index State: Since CRACK is a lazy operation, the index can exist in two concurrent states, the true state and a dynamic state.
+Index State: Since CRACK is a lazy operation, the index has two concurrent states, the true current state and a dynamic buffered state.
 The true state reflects the current structure of the index, while the dynamic one tracks its expected future structure for when $CRACK$ is executed, which synchronizes the two states.
-Maintaining both is essential.
-$REFINE$ operations directly alter the current index state, while cracking decisions are made based on the dynamic one, leading to scenarios that require us to undo buffered crack decisions.
-Initially, the heuristic rules considered some crack candidates as beneficial, and CrackIVF buffered them, but they have since been altered.
-Rolling back the dynamic state avoids bad cracks from being committed and having a situation where we commit decisions to the index that we know to be suboptimal (Line 21 Algorithm 1).
-To track the dynamic, uncommitted, state of the index, we define:
-(1) $\text{State}_{\text{dyn}}=(A_{P}^{\text{dyn}},D_{P}^{\text{dyn}},H_{C}^{\text{dyn}})$
-which consists of dynamic assignments $A_{P}^{\text{dyn}}$, dynamic point distances $D_{P}^{\text{dyn}}$, and a histogram $H_{C}^{\text{dyn}}$ of true and buffered crack sizes.
-Similarly, the true state is defined as:
-(2) $\text{State}_{\text{true}}=(A_{P}^{\text{true}},D_{P}^{\text{true}},H_{C}^{\text{true}})$
-representing the true, committed state of assignments, distances, and crack sizes in the index.
-The initialization of the above states happens during the index setup time.
-More specifically, in our FAISS-based implementation, during an “.add()“ operation.
-This is when the data vectors are assigned to their nearest partitions by calculating the distances to each partition representative.
-We extend this operation to return to the assignments and distances computed, which were an intermediate result.
-The state held in $A_{P}^{\text{dyn}}$,$D_{P}^{\text{dyn}}$ and $H_{C}^{\text{dyn}}$, is only consistent with the actual state of the index after $CRACK$ operations and before buffering any cracks.
+$REFINE$ operations directly alter the current index state, while cracking decisions are made based on the dynamic state.
+This can lead to scenarios that require us to undo buffering decisions.
+This is done to avoid bad cracks from being committed and having a situation where we commit decisions to the index that we know to be suboptimal (Line 21, Algorithm 1).
+To track the index state, we define:
+(1) $\text{State}_{\ast}=(A_{P}^{\ast},D_{P}^{\ast},H_{C}^{\ast})$
+where $\ast\in\{\text{dyn},\text{true}\}$ denotes either the dynamic uncommitted or true committed state.
+The dynamic state consists of dynamic assignments $A_{P}^{\text{dyn}}$, dynamic point distances $D_{P}^{\text{dyn}}$, and a histogram $H_{C}^{\text{dyn}}$ of both the true and buffered crack sizes.
+Similarly, the true state represents the true, committed state of assignments, distances, and crack sizes in the index.
+We extend the FAISS-IVF “.add()” operation to return the assignments and distances computed, which are an intermediate result, in order to initialize the state tracking.
+The state held in $A_{P}^{\text{dyn}}$$D_{P}^{\text{dyn}}$ and $H_{C}^{\text{dyn}}$, is only consistent with the actual state of the index after $CRACK$ operations and before buffering any cracks.
 If no cracks are buffered and only $REFINE$ operations are executed, the dynamic and true state are always consistent.
-All of the above state metadata is used after the SEARCH operation, along with the search results $I,D,C_{local}$ to classify each query as a good or bad crack candidate (Section 4.2).
+All of the above state metadata is used after the SEARCH operation, along with the search results $I,D,C_{visited}$ to classify each query as a good or bad crack candidate (Section 4.2).
 Rollback of buffered state: Consider the following scenario.
-A $CRACK$ operation was just executed so at $t_{0}$ $State_{true}=State_{dyn}$.
-After a few queries, at $t_{1}$, assume that 2 good crack candidates have been buffered, both in the $C_{local}$ region of the vector space and thus now $State_{true}\neq State_{dyn}$.
-At a later point, $t_{2}$, a $REFINE$ is executed, also in $C_{local}$ refining it to $C^{\prime}_{local}$ and changing $S_{true}$.
-Since $REFINE$ operates to improve the current state eagerly, the decision that the buffered cracks are good was taken prior, at $t_{1}$ based on $C_{local}$, and thus they may now have had their points stolen back, and assigned to the refined cracks in $C^{\prime}_{local}$.
-Thus even though still $|C_{buffered}|=2$, it may now contain bad cracks.
-So when at $t_{3}$, it is a good time to execute a $CRACK$ operation again, all buffered cracks are re-evaluated, and only the good cracks that remain are committed, ensuring that in the end, we are back to $State_{true}=State_{dyn}$.
-This logic can be seen in lines 21-25 of Algorithm 1.
-Memory Overhead and Optimizations: The additional memory cost of maintaining $State_{\text{dyn}}$ and $State_{\text{true}}$ is $O(4|P|+2|C|)$, which is negligible since $d\gg 4$, so in our in-memory implementation that data alone are $O(|P|*d)$.
+A $CRACK$ operation was just executed at time $t_{0}$, so $State_{true}=State_{dyn}$.
+After a few queries, at $t_{1}$, assume that 2 good crack candidates have been buffered, both in the same $C_{local}$ region of the vector space and thus now $State_{true}\neq State_{dyn}$.
+At a later point, $t_{2}$, a $REFINE$ is executed, also in $C_{local}$ changing the assignments to $C^{\prime}_{local}$, affecting $S_{true}$.
+Since $REFINE$ improves the current state eagerly, the decision that the buffered cracks are good was taken at a prior time $t_{1}$, on a prior state $C_{local}$.
+Thus, at $t_{2}$, during $REFINE$ the buffered cracks may now have had their points stolen back, and assigned to other refined cracks in $C^{\prime}_{local}$.
+Now, while still $|C_{buffered}|=2$, it may now contain bad cracks.
+Finally, when at $t_{3}$, it is a good time to execute a $CRACK$ operation again, all buffered cracks are re-evaluated, as shown in lines 22-24 Algorithm 1, and only the good cracks that remain are committed.
+Memory Overhead and Optimizations: The additional memory cost of maintaining $State_{\text{dyn}}$ and $State_{\text{true}}$ is $O(4|P|+2|C|)$.
+It is negligible since $d\gg 4$ and the in-memory data alone are $O(|P|\times d)$.
 Additionally, these array structures are freed once the index converges or incoming queries stop.
 In our case, it enables an efficient compute-storage trade-off, deferring cracking costs instead of paying them per query.
-There are a number of optimizations that a fully fledged implementation can make use of.
+There are many optimizations that a full-fledged implementation could utilize.
 For embedding lakes, where vectors are stored on disk, state tracking arrays can be stored alongside data points instead of in the index (irrelevant for our in-memory FAISS implementation).
-Since queries for an individual index are typically infrequent, $State_{\text{dyn}}$ can be committed or flushed between queries, while $State_{\text{true}}$ is recomputed as needed.
-Full memory overhead is only necessary during bursty query loads, where cracking must be buffered to reduce cumulative overhead and prevent long tail response times.
+Since queries for an individual index are typically infrequent, $State_{\text{dyn}}$ can be committed or flushed between queries, while $State_{\text{true}}$ recomputed as needed.
+Full memory overhead is only necessary during bursty query loads with limited hardware resources, where cracking must be buffered to reduce cumulative overhead and prevent long tail response times.
 Finally, the cracking overhead can be hidden from query response times since results are available before any cracking operations begin.
 A separate thread pool can handle cracking operations while queries use the current index state, provided $CRACK$ and $REFINE$ remain atomic to prevent inconsistencies.
-Prior work on the related problem of index maintenance has demonstrated similar approaches to minimizing the overhead of mutating the ANN index state (Xu et al., 2023) with concurrent operations, we further cover these approaches in Section 6.
-Our work assumes a continuous query load at the system’s capacity, leveraging inter-query parallelism.
-As a result, the memory overhead for index state tracking is unavoidable in our implementation if we aim to maintain high performance.
+Prior work on the related problem of index maintenance (Xu et al., 2023) demonstrated how to concurrently execute queries by dispatching such tasks to background threads.
+In our work, the entire thread pool is dedicated either to search or cracking operations.
+We also assume a continuous query load at system capacity.
+As a result, the memory overhead for index state tracking is unavoidable.
+We further cover index maintenance prior work in Section 6.
 4.2. Where to apply build operations?
-There is a constant stream of crack and refinement candidates, and as observed in Section 3, not all regions of the vector space are equally important, and not all operations applied lead to improvement of index performance.
-Thus, a set of rules is needed to distinguish good candidates from bad ones.
-The heuristic rules provided in this section aim to distill the learning from our initial observations to capture independent cases where we want to add a new crack or refine a region.
-At this stage, they are not an exhaustive list, as we plan to extend them in future work.
-Where to CRACK: The CrackHeuristic determines whether an incoming query is a good or bad candidate for cracking, as shown in Algorithm 1.
-It evaluates the stolen point IDs $I_{\text{steal}}$ and the dynamic state $State_{\text{dyn}}$ to return a True/False decision.
-If classified as good, the query is buffered for a future $CRACK$.
-By default, we aim to expand the index size, as our insights suggest that optimal performance is typically achieved at some maximum number of partitions.
-Since $REFINE$ does not add partitions, we assume every incoming query is a good candidate to have a constant stream of candidates that can expand the size of a CrackIVF index and only reject them if they pass any of our heuristic rules.
-CrackHeuristic: A query is considered a bad candidate (returning False) if any of the following conditions hold:
-Don’t Crack: Rule 1 rejects queries with too few stolen points:
+Each arriving query is a crack candidate, and the visited region during index traversal is a candidate for refinement.
+This results in a continuous stream of candidates of varying quality.
+As discussed, not all vector space regions are equally important, and not all operations lead to performance gains.
+The goal is to maintain a stream of high-quality candidates, classified from incoming queries, enabling the index to grow by adding cracks in frequently accessed regions that need further partitioning, and refining poorly clustered regions with K-means.
+CrackIVF employs two heuristics, acting as a binary classification model for candidates.
+Each uses a set of decision rules that are distilled from our empirical observations of IVF index behavior.
+We do not claim to have defined an exhaustive list of decision rules for the model.
+Changing or learning the model and the parameters while still achieving the stated goal above, would not change the overall design of the index (Algorithm 1).
+Where to CRACK (CrackHeuristic): We observe that the optimal performance is achieved at some maximum number of partitions (Figure 2(a)).
+Thus, for it to grow, we treat each incoming query as a good crack candidate by default and define a set of decision rules to classify when a query is a bad candidate.
+Query candidates classified as good, are buffered until a future $CRACK$ operation.
+A bad candidate (returning False) is if any of the following hold:
+Don’t Crack: Rule 1 rejects queries that steal too few points:
+(2) $|I_{\text{steal}}|<\text{min\_pts}$
+Don’t Crack: Rule 2 prevents excessive partitioning, constraining local partition count based on the local number of points:
 […]
-Don’t Crack: Rule 2 prevents excessive partitioning, constraining the number of partitions within a region relative to the number of points:
-[…]
-[…]
-Where to REFINE: The RefineHeuristic determines whether a local region in the index is a good candidate for refinement, returning a $True/False$ value for if a refinement should take place.
-Unlike $CRACK$, where the latency overhead is amortized by buffering, $REFINE$ is applied eagerly and is also an expensive operation, running K-Means locally, calculating new point assignments, and reorganizing the physical storage.
-Additionally, if the index continues to grow in the future, any previously refined region may need further refinement, making premature refinements inefficient.
-It is much better to relax rules for when to refine when the index has reached its maximum size and has also accumulated a larger cost budget.
-Therefore, we do not want to be calling $REFINE$ repeatedly, unless necessary, and our default behavior is to assume that every incoming query local region is NOT a good candidate for refinement unless it satisfies specific heuristic rules that indicate a significant local partition imbalance.
-Imbalance affects tail response time negatively and increases the variance since even though queries may be accessing the same number of partitions, they may be scanning a vastly different number of points.
-Additionally, running K-means improves the centroid location of the crack representatives, making them better fit the underlying data distribution, which also has downstream benefits (Mohoney et al., 2024).
-Various variance measures exist for detecting imbalances in cluster assignments (Hastie et al., 2009).
-We selected the following subset as they capture different aspects of imbalance, both locally and globally.
-RefineHeuristic: A local region is considered a good candidate for refinement (returning True) if any of the following conditions hold:
-Refine: Rule 1 captures local imbalance using the coefficient of variation:
-[…]
+We set $\text{min\_pts}=2$ in order to avoid adding empty cracks, and $\text{pts\_crack\_thr}=64$ to ensure an average of at least 64 points per crack locally.
+Rule 2 is especially important in high skew datasets, where most crack candidates are for the same region.
+Where to REFINE (RefineHeuristic): We observe that refining an index improves performance at each index size (Figure 2(a)), by better distributing points across partitions and repositioning centroids to better fit the data.
+Optimal performance is reached at the maximum index size, so we want to avoid getting stuck refining a small index.
+Unlike $CRACK$, where latency is amortized by buffering, $REFINE$ is eagerly applied to immediately improve the current index, but it is also a costly operation.
+Thus, we treat each incoming refine candidate as bad, leaving more of the cost budget for cracks so that the index can grow.
+A refine candidate is the local visited region of the query, visited during search.
+It is a good candidate for refinement (returning True) if any of the following hold:
+Refine: Rule 1 captures local imbalance:
+(4) $\text{CV}_{\text{local}}=\frac{\sigma_{\text{local}}}{\mu_{\text{local}}}>\text{cv\_max}$
 where $\sigma_{\text{local}}$ and $\mu_{\text{local}}$ are the standard deviation and mean of cluster sizes in the local region.
 A high coefficient of variation indicates substantial local imbalance, warranting refinement.
-Refine: Rule 2 captures extreme values using local spread:
-[…]
-This ensures that refinement is triggered when large disparities exist between the smallest and largest clusters within a local region.
-Refine: Rule 3 captures imbalance based on the global distribution:
-[…]
-[…]
+Refine: Rule 2 captures imbalance using the global distribution:
+(5) $\exists c_{1},c_{2}\quad\text{s.t.}\quad S(c_{1})\leq{\text{cutoff}_{\text{low}}},$
+$S(c_{2})\geq{\text{cutoff}_{\text{high}}}$
 where $c_{1},c_{2}$ are cracks in the local region, and $S(c)=H_{C}^{\text{true}}[c]$ represents the size (i.e., number of assignments) of crack $c$.
-[…]
-These thresholds ensure that refinement is applied only when clear imbalance exists, maintaining index efficiency.
-An optimized implementation could relax these rules when more computational budget is available, allowing refinement of less imbalanced regions.
+The $\text{cutoff}_{\text{low}}$ and $\text{cutoff}_{\text{high}}$ correspond to cluster sizes at the size_prctlsize\_prctl and 100size_prctl100-size\_prctl percentiles of the global partition sizes.
+This rule captures the unlikely scenario that a large and small crack globally are next to each other locally, indicating points can be better distributed.
+We set $\text{cv\_max}=2$, meaning refinement is triggered if cluster sizes vary by at least twice the local mean, and $\text{size\_prctl}=10$, defining small and large cracks based on the 10th and 90th percentiles of global cluster sizes.
 4.3. When to apply build operations?
-To mitigate the overhead of constant physical reorganization, which is expensive in high-dimensional IVF indexes, we use a build budget that restricts the fraction of time dedicated to $CRACK$ and $REFINE$ operations.
-We define $T_{\text{build}}$ as the total measured time spent on previous build operations, which include historical $CRACK$ and $REFINE$.
+To mitigate the overhead of constant physical reorganization, we use a build budget to restrict the fraction of time dedicated to $CRACK$ and $REFINE$ operations.
+We define $T_{\text{build}}$ as the total measured time spent on previous build operations.
 Similarly, $T_{\text{search}}$ represents the cumulative time spent on all past search operations.
-When considering a potential build operation, we estimate its execution cost using $\hat{T}_{f}$, which captures the expected time required to perform either a $CRACK$ operation under the current dynamic state ($State_{dyn}$) or a $REFINE$ operation on the local region of the current refine candidate.
-Here, $f\in\{\textsc{CRACK},\textsc{REFINE}\}$ specifies the build operation considered.
-To enforce the budget constraint, we ensure that the total time spent on past and projected build operations remains within a fraction $\alpha$ of the system’s total runtime:
+When considering a potential build operation, we estimate its execution cost using $\hat{T}_{f}$, which captures the expected time required to perform a $CRACK$ operation in the current dynamic state or a $REFINE$ operation in the true state.
+The build operation considered is specified by $f\in\{\textsc{CRACK},\textsc{REFINE}\}$.
+To enforce the budget constraint, the total time spent on past build operations in addition to the estimated time if the current one is executed must remain within a fraction $\alpha$ of the system’s total runtime:
 (6) $T_{\text{build}}+\hat{T}_{f}\leq\alpha\cdot(T_{\text{build}}+T_{\text{search}}+\hat{T}_{f}).$
-$T_{\text{build}}$ and $T_{\text{search}}$ are directly measured from execution, while $\hat{T}_{f}$ is derived from a predictive cost model, which is explained in the following subsection.
-The parameter $\alpha$ governs the proportion of total execution time that can be allocated to indexing operations, ensuring a balanced trade-off between indexing efficiency and query performance.
-This approach allows for a consistent and adaptive mechanism to regulate both cracking and refining decisions under varying workload conditions.
-Varying $\alpha$ is a way of expressing an expectation about the number of queries that will arrive.
-A large $\alpha$ allocates a larger build budget, which is useful if we are optimistic about receiving a large number of queries in the future, thus spending more time to build an index is beneficial.
-A smaller $\alpha$ restricts the build budget, indicating that we are pessimistic and do not expect many more queries to arrive.
-To have a balance, in all our experiments, we set $\alpha=0.5$, to ensure a maximum of 50% of the total time at any given point is spent on build operations.
-This budget ensures that the build operation overhead does not dominate the overall execution time.
+$T_{\text{build}}$ and $T_{\text{search}}$ are directly measured from the execution time, while $\hat{T}_{f}$ is estimated using a predictive cost model.
+The parameter $\alpha$ defines the proportion of total execution time that can be allocated to indexing operations.
+We set $\alpha=0.5$, which ensures a maximum of 50% of the total time at any given point is spent on build operations.
 As more queries arrive, CrackIVF will progressively have a larger budget to apply for physical reorganization operations.
@@
 Finally, there is a minimum size constraint on the number of buffered cracks before a $CRACK$ is executed.
-Specifically, in our implementation, we set that minimum to 20% of the current index size, $|C_{buffered}|>C_{min}$, where $C_{min}=0.2*|C|$ (Line 19 Algorithm 1).
-This helps avoid situations where a $REFINE$ operation applied on the current index achieves similar performance improvements, while allowing us to keeping the current index size without increasing it.
-Predictive Cost Model
-The cost estimation of both CRACK and REFINE operations happens after every incoming query.
-Thus, it is crucial to maintain low latency when making a prediction while at the same time maintaining an accurate enough estimate to satisfy our usage needs.
-To achieve this balance, we employ a simple first-order linear model and fit it using multivariate regression (Jain, 1990).
+The idea comes from Figure 2(a), where if a REFINE can achieve similar performance improvement at the current index size, it is better to wait until more cracks have been buffered, so that the performance improvement from index growth is substantial.
+Specifically, in our implementation, we set that minimum to 20% of the current index size, $C_{min}=0.2\times|C|$ (Line 19 Algorithm 1).
+Predictive Cost Model: Estimating the execution time of the CRACK and REFINE operations happens after every incoming query.
+Thus, we must maintain low latency when making a prediction, while have accurate enough cost estimates to correctly use the budget constraint Eq.
+6.
+For this, we use a simple first-order linear model fitted using multivariate regression (Jain, 1991).
 Each build function $f$ comprises a sequence of procedures, or kernels, indexed by $i$ and executed sequentially, as shown in Algorithm 2 and Algorithm 3.
-Among these, updating the index state is negligible, and the following kernels are the dominant factors of each function’s execution time:
+The execution time is dominated by the following kernels:
 CRACK: (i) Get Local Region, (ii) Commit Reorg (w/ $A_{P}^{\text{dyn}}$), (iii) Update Centroids.
 REFINE: (i) Get Local Region, (ii) Local k-Means, (iii) Commit Reorg (w/o $A_{P}^{\text{dyn}}$).
-Since the above kernels in our implementation execute sequentially without overlap, the total execution time $T_{f}$ is simply the sum of the execution times of its kernels.
 (7) $T_{f}=\sum_{i}T_{f,i},\quad\forall f\in\{\textit{CRACK},\textit{REFINE}\}.$
+These kernels are executed sequentially so the total execution time $T_{f}$ is simply the sum of kernel execution times.
 Each kernel $i$ contributes a latency $T_{f,i}$, which we model as a linear function of its two dominant cost factors: computational complexity and data movement.
-Specifically, for each kernel $i$ in function $f$, we train the following simple predictive model:
+For each, we learn the following predictive model:
 (8) $T_{f,i}=w_{f,i}^{(1)}\cdot C_{f,i}+w_{f,i}^{(2)}\cdot D_{f,i}+b_{f,i},$
 where $C_{f,i}$ represents the dominant computational complexity of kernel $i$, and $D_{f,i}$ denotes the dominant data movement cost.
-The terms $w_{f,i}^{(1)}$ and $w_{f,i}^{(2)}$ are learned coefficients corresponding to computation and data movement, respectively.
-Finally, $b_{f,i}$ is a learned parameter that accounts for fixed kernel execution overheads.
+The terms $w_{f,i}^{(1)}$ and $w_{f,i}^{(2)}$ are learned coefficients for computation and data movement and $b_{f,i}$ is a learned parameter that accounts for fixed overheads.
 We gather measurement with micro-benchmarks for each kernel, by varying input parameters and measurig execution latency to fit a regression model, learning $w_{f,i}^{(1)},w_{f,i}^{(2)},b_{f,i}$.
-Our approach eliminates the need for handcrafted analytical performance models that map the kernels into the number of cycles per operation and are tied to specific implementations and hardware configurations, making them hard to adapt or extend.
-With our method, switching hardware or optimizing kernels only requires re-running microbenchmarks and refitting the linear model.
-Another advantage is the low computational cost of inference, enabling continuous use during query execution.
-However, modeling execution time as a simple linear function has limitations.
-While it captures dominant compute and memory trends, it overlooks system-level nonlinearities.
-Factors like caching effects, NUMA constraints, SIMD optimizations (e.g., in FAISS), and thread contention introduce discrete performance variations that a linear model cannot fully express.
+Table 2. Features used for each kernel. For REFINE, $C_{\text{buffered}}=1$
 Kernel Compute Data Movement
-Get Local Region* $|C_{\text{buffered}}|\times|C|$ $|P_{\text{local}}|\times d$
+Get Local Region $|C_{\text{buffered}}|\times|C|$ $|P_{\text{local}}|\times d$
 Commit Reorg (w/o $A_{P}^{\text{dyn}}$) $|P_{\text{local}}|\times|C_{\text{local}}|\times d+|C|$ $|P_{\text{local}}|\times d$
@@
 Local K-Means $n_{\text{iter}}\times|P_{train}|\times|C_{\text{local}}|\times d$ $|C|\times d$
-Table 2. Features we used for each kernel. * For REFINE, $C_{\text{buffered}}=1$
-Nevertheless, our empirical results show that the cost model performs well across all benchmarks.
-Importantly, we always use the true measured times $T_{\text{build}}$ and $T_{\text{search}}$ in our budget constraint Equation 8.
-This ensuring any overestimation by the predictive model is corrected after execution, and an underestimation will hold any future build operations until sufficient queries arrive, increasing the total budget.
-In our testing, the model achieved moderate to high $R^{2}$, explaining between 40 and 95% of the total variation, with the $RMSE$ remaining in the millisecond to low second range.
-We find this to be acceptable performance from our model since we benefit from low inference latency, and the cost budget-based control mechanism works effectively on all datasets, as we show in our experiments.
+Our approach removes the need for exact analytical models coupled to specific implementations and hardware.
+Switching hardware requires re-running micro-benchmarks and refitting the linear model.
+The main limitation is that execution time is modeled as a linear function.
+While this captures dominant compute and memory trends, it overlooks system-level nonlinearities such as caching effects, NUMA constraints, and SIMD optimizations.
+Despite this, our empirical results show that the cost model performs well across all benchmarks.
+Importantly, we always use the true measured times $T_{\text{build}}$ and $T_{\text{search}}$ in the budget constraint (Equation 6), which help correct over and under estimations.
+In our tests, the model achieved moderate to high $R^{2}$ scores, explaining 40–95% of the total variance, with RMSE ranging from milliseconds to a few seconds.
 5. Experiments
-5.1. Experimental Setup
-We run on a dual-socket AMD EPYC 7V13 system (128 cores, 3.7 GHz, 512GB RAM) from the AMD-ETHZ HACC cluster (Moya et al., 2023a, b, 2024).
-Unless otherwise stated, experiments use 16 cores, the number at which the memory bandwidth saturates, and use an equal query batch size to benefit from inter-query parallelism.
-Table 3. AV-Tree vs. CrackIVF
-Cumulative Time at i-th Query (seconds)
-$10^{0}$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ (Total)
-AV-Tree 0.078 0.747 7.601 75.369 825.709 1264.175
-CrackIVF 1.107 1.166 1.738 15.983 98.295 516.881
-Datasets: We evaluate on standard ANN search benchmarks for text and image-based retrieval, GloVe(Pennington et al., 2014), SIFT(Jégou et al., 2011), DEEP(Babenko and Lempitsky, 2016), and Last.fm(Bertin-Mahieux et al., 2011).
-The latter, being an example with a high skew, is used for music recommendation.
-The SIFT1M and SIFT10M, contain 128-dimensional SIFT descriptors, while DEEP10M consists of 96-dimensional embeddings from a fully connected GoogLeNet model (Szegedy et al., 2015).
-These datasets are subsets of the original billion-scale collections, with 1M and 10M referring to the number of data points in each slice.
-Both use L2 distance as a metric.
-GloVe, sourced from ANN-Benchmarks (Aumüller et al., 2020), represents text retrieval.
-It contains 1.18M varying dimensional data points and uses cosine similarity, which we implement by l2-normalizing and computing inner product (IP) similarity.
-Last.fm is also taken from (Aumüller et al., 2020); it contains 64-dimensional vectors and uses the IP similarity metric.
-We found that the provided query set is highly skewed, allowing us to test how CrackIVF behaves in such a scenario.
-Last.fm provides 50k unique queries, while the other datasets provide 10K unique queries.
-Thus, we replicate them until we match our target scales (up to 1M).
-Baselines: Our comparisons cover both exact nearest neighbor (ENN) and approximate nearest neighbor (ANN) search approaches.
-For ENN, we use Brute Force, i.e., a linear scan of the data and AV-Tree (Lampropoulos et al., 2023), as a cracking baseline.
-For ANN search, we compare against FAISS IVF indexes, which in our figures are indicated as IVFFlat, with the number of partitions chosen in parentheses, e.g., IVFFlat (1000).
-We vary the number of partitions from 1,000 to 16,000, following FAISS guidelines (Facebook Research, 2025), which recommend this range for our dataset scales.
-Our method, CrackIVF is initialized with 100 partitions to maintain near brute force startup time.
-All experiments have the same fixed parameters, with values set as explained in (Sections 4.2, 4.3).
-The $\alpha$ parameter to control the budget remains fixed to $0.5$.
-Allowing, a build budget equal to 50% of the total time.
-Finally, for the partition-based indices, we have CrackIVF and IVFFlat baselines, where the $nprobe$ is set to always achieve 90–95% Recall@10.
-The QPS-Recall plots show the Pareto frontier when varying $nprobe$ across all indexes.
 (a) SIFT1M: QPS vs. Recall
@@
 Figure 3. Queries Per Second (QPS) vs. Recall and Time per each Query batch for the entire trace across different datasets.
+We run on a dual-socket AMD EPYC 7V13 system (128 cores, 3.7 GHz, 512GB RAM).
+All experiments, except for Table 3, use 16 cores and a batch size of 16, the point at which memory bandwidth saturates.
+IVF and CrackIVF share the same SEARCH operation in FAISS and benefit equally from inter-query parallelism.
+We evaluate on standard ANN search benchmarks, GloVe(Pennington et al., 2014), SIFT(Jégou et al., 2011), DEEP(Babenko and Lempitsky, 2016), and Last.fm(Bertin-Mahieux et al., 2011).
+SIFT1M and SIFT10M contain 128-dimensional SIFT descriptors; DEEP10M consists of 96-dimensional embeddings, with 1M and 10M being the number of data points in each slice.
+Both use L2 distance as a metric.
+GloVe contains 1.18M varying dimensional embeddings and uses cosine similarity, which we implement by l2-normalizing and computing the inner product (IP).
+Last.fm, contains 64-dimensional vectors and IP metric.
+The Last.fm query set is highly skewed, allowing us to test CrackIVF in such cases.
+Last.fm provides 50k unique queries, while the other datasets provide 10K unique queries.
+We replicate them until we match our target scales (up to 1M).
+For ENN, we compare against Brute Force, and AV-Tree (Lampropoulos et al., 2023), which is a cracking baseline.
+For ANN search, we compare against FAISS IVFFlat indexes with the number of partitions chosen shown in parentheses, e.g., IVFFlat (1000).
+We vary the number of partitions following FAISS guidelines (Facebook Research, 2025), for our dataset scales.
+For CrackIVF, we use fixed parameters values for the control mechanisms, set as explained in (Sections 4.2, 4.3).
+We consider the index converged when the crack decision rules stop triggering and refines keep targeting the same region without effect, thus the index can no longer grow in size, nor improve at the current size.
+For CrackIVF and IVFFlat, nprobenprobe is set to always achieve 90–95% Recall@10 in ANN.
+CrackIVF is initialized with 100 partitions on all experiments.
+We set this parameter by choosing the largest possible starting partitions that still give minimal startup cost.
+Because of FAISS parallelism, at 16 threads, the startup time for 100 partitions is identical or faster than starting with 1 partition.
+On a single thread for Table 3, the difference between them is only $\approx$120 ms.
 5.1. Comparison with AV-Tree:
-To compare with AV-Tree (Lampropoulos et al., 2023) fairly, we set our CrackIVF to run with a single thread and make sure we probe enough partitions to get an average of more than $99\%$ recall across all query scales.
+Table 3. AV-Tree vs. CrackIVF. Cracking-based baselines.
+Cumulative Time at i-th Query (seconds)
+$10^{0}$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
+IVF-1K 12.57 12.59 12.76 14.48 32.10 205.40 1938.60
+IVF-5K 178.17 178.18 178.35 179.93 195.94 354.63 1943.02
+AV-Tree 0.08 0.75 7.48 75.43 830.0 1274.85 5721.29
+CrackIVF 1.10 1.16 1.73 13.18 55.33 233.11 1531.40
+To the best of our knowledge, AV-tree (Lampropoulos et al., 2023) is the only similarity search index, whose underlying idea is based on cracking.
+It is a tree-based ENN cracking index targeting short-lived data up to one thousand queries.
+Unlike AV-Tree, CrackIVF is an IVF-based ANN cracking index, implemented around FAISS, offering parallelism, and targeting high dimensional RAG applications which could reach much larger query volumes.
+As AV-Tree is single-threaded and an ENN index, to compare fairly, we set CrackIVF and the IVF baselines to run with a single thread and make sure we probe enough partitions to get an average of more than $99\%$ recall across all query scales.
 We set the cracking threshold of AV-Tree to 128, which we find gives the best total runtime.
-We show the cumulative time for a trace of $100$K queries over the SIFT 1M dataset in Table 3.
-Unlike ours, AV-Tree does not pre-build the index, therefore, it has an advantage over us for the first 10 queries.
-However, our method processes each query faster and outperforms AV-Tree by a large margin after the first few queries.
-The cumulative time of AV-Tree increases at a slower rate with more queries, meaning that the query time is getting faster, and AV-Tree is improving over time.
-Nevertheless, our CrackIVF is 2.45x faster than AV-Tree in the end, even on a single-threaded ENN search.
-This shows the effectiveness of ANN approaches, where a less than 0.01% Recall error leads to a large gain in performance.
-For the rest of the experiments, we focus on the ANN search setting.
+We show the cumulative time for a trace of 11M queries over the SIFT 1M dataset in Table 3.
+At the start, the $10^{0}$-th query time includes the index construction costs of CrackIVF and IVF baselines.
+CrackIVF is initialized as a small 100 partition index which grows as more queries arrive, whereas AV-Tree does not pre-build any index.
+In this experiment, AV-Tree has an advantage over CrackIVF for the first 10 queries, as it can start answering immediately.
+However, CrackIVF processes each query faster and outperforms AV-Tree by a large margin after the first few queries.
+The cumulative time of AV-Tree increases at a slower rate with more queries, showing AV-Tree is improving over time.
+Our CrackIVF is 3.7x faster than AV-Tree in the end of the trace.
+As cracking baselines, both CrackIVF and AV-Tree, vastly outperform IVF indexes on the low query scale.
+AV-Tree continues to outperform up to 1000 queries, matching original claims (Lampropoulos et al., 2023).
+CrackIVF performance benefits continue even at 1 million queries.
+In the following experiments we focus on ANN.
 5.2. Does CrackIVF improve over time?
-To measure how CrackIVF improves over time, we provide two graphs for each of the datasets and group them in Figure 3.
-The QPS vs.
-Recall plots measure the maximum achievable Queries Per Second (QPS) across varying Recall targets.
-This is the trade-off under which the ANN search operates.
-Having a large gain in search speed (QPS) for a small decrease in search quality (Recall).
-By the end of the 1 million query trace, indicated by CrackIVF-END (orange line), we are consistently at the Pareto frontier of QPS-Recall across on all datasets.
+We measure how CrackIVF improves over time in Figure 3.
+QPS vs.
+Recall plots show the search performance trade-off in which ANN search operates under.
+In our experiments, by the end of the query trace, CrackIVF-END is consistently at or near the Pareto frontier across all tested datasets.
 On SIFT10M and Last.fm, CrackIVF is by far the best-performing index.
-Specifically on Last.fm, CrackIVF with 1255 final partitions, achieves 50% higher QPS for Recall@10 = 0.9, than a 10x larger IVFFlat(16000) index.
-We attribute this to CrackIVF’s approach of only allocating new cracks based on the query distribution, and since Last.fm is the dataset with the highest skew, where only  30% of the vector spaces is accessed (Figure 2(b)), we can fully partition that  30% of the space with a relatively small number of cracks to achieve the best performance.
-In Figure 3, we also plot the Response Time per query batch across the entire 1 million query trace.
-We can see that CrackIVF consistently improves over time, showing drops in response time after incremental reorganization operations.
-The outliers can all be attributed to these $CRACK$ and $REFINE$ reorganization operations.
-This is evidence for having control mechanisms to ensure expensive operations are only applied when necessary and why CrackIVF relies on buffering crack operations.
-The downstream of buffering and committing a large number of multiple cracks at once is only visible in one dataset, Last.dm, where it creates a step-like response in the data, as seen in Figure 3(f).
-This sudden improvement in performance when adding a lot of cracks can also be seen in Figure 2(a), when going from 1000 to 5000 partitions.
+Specifically on Last.fm, CrackIVF with 1264 final partitions, achieves $\approx$50% higher QPS for Recall@10 = 0.9, than any other index.
+We attribute this to CrackIVF’s approach of allocating new cracks following the query distribution.
+Since Last.fm is the dataset with the highest skew (Figure 2(b)), we can fully partition the accessed space with a small number of cracks, achieving the best performance.
+In Figure 3, we also plot the response time per query batch across the entire query trace.
+CrackIVF search performance improves over time after incremental reorganization operations.
+The vast majority of outliers are due to the individual $CRACK$ and $REFINE$ reorganization operations.
+In our experiments, $CRACK$ operations are less than 0.03% and $REFINE$ less than 0.6% of the query trace.
 5.3. Does CrackIVF Minimize Cumulative Time?
-Our core motivation in this work has been to provide an index that can be deployed across data lakes under scenarios where the number of queries that will arrive for each dataset and utilize the index is unknown.
-The ideal index for such a scenario should be able to minimize the cumulative time spent on both build and search operations in cases where a dataset may receive a few queries as well as millions.
+We run a trace of 1 million queries for each dataset and include the upfront build cost for all baselines as the time before they answer the first query (Figure 4).
+The smaller 1000 partition IVF indexes have relatively low startup costs but do not scale well past $\approx$10k queries.
+Larger indexes have a huge upfront build cost, which only pays off if a dataset receives ¿1 million queries.
+Brute-force has the minimum initialization time, making it the better choice when the number of queries is low (¡100).
+In our experiments, CrackIVF achieves near-minimum cumulative time across all query scales.
+It achieves almost 1000x lower initialization time than larger indexes (Figure 4(d)) and can answer 100K to 1M queries before comparable indexes finish building.
+For example on SIFT10M (Figure 4(b)), the initialization time cost is $\approx$100x lower than IndexFlat(5000) and almost 1000x lower than IndexFlat(16000), yet CrackIVF answers 100k and 1M queries respectively, before the indexes finish building.
 (a) SIFT1M
@@
 Figure 4. Cumulative time plots across datasets.
-We evaluate this specific aspect of index performance, across all datasets, with the cumulative time experiments, shown in Figure 4.
-For our measurements, we run a trace of 1 million queries for each dataset.
-In all baseline methods, we include the upfront build cost as the time before they answer the first query.
-The smaller IVFFlat (1000) indexes have relatively low startup costs but do not scale well past  10k queries since their cumulative time rises sharply due to having higher query batch response times than the alternative.
-The larger indexes, beyond 5000 partitions, have a huge upfront build cost, which may pay off but only if a dataset receives ¿1 million queries.
-The “no index“ approach of the Brute-force method has the minimum initialization time and is thus the best choice when the number of queries is low (¡100).
-CrackIVF consistently achieves near-minimum cumulative time across all query scales.
-It achieves up to 1000x lower initialization time than larger indexes (Figure 4(d)) and can consistently answer 100K to 1M queries before comparable indexes finish building.
-For example on SIFT10M (Figure 4(b)), the initialization time cost is 10x lower than IndexFlat(5000) and 100x lower than IndexFlat(16000), yet it processes 100K and 1M queries respectively, before they baselines indexes complete construction, while at the same time converging to higher QPS-Recall performance.
-Similar results can be observed in all other datasets.
-The fact that CrackIVF achieves the Pareto optimal result in QPS-Recall by the end of our trace means that, by extrapolation, it will always maintain the minimum cumulative time across all query scales beyond 1M queries.
+5.4. Does CrackIVF Reduce Total Distance Computations for Build Operations?
+IVF indexes incur full indexing cost upfront over the entire vector space.
+Indexing large datasets (100M–1B+) in reasonable time requires GPUs, but even then, K-means training can take hours due to the increasing number of training points and partitions, which have a multiplicative effect in the number of pairwise distance computations (see Table 2 K-means Compute).
+CrackIVF amortizes this cost over time using $REFINE$ operations, which apply K-means locally on small subsets of points and partitions.
+As a result, total distance computations are spread across time and space and can potentially be handled on cheaper hardware (e.g., CPUs).
+Figure 5 shows the cumulative distance computations during index build, excluding search operations.
+CrackIVF pays this cost gradually, and after the index converges, the final distance computations can be orders of magnitude lower than larger IVF indexes (Figure 5(a)5(f)).
+Despite having comparable or better performance (Figure 3(a)3(k)).
+Convergence occurs when the arrival of new crack candidates stops and refines do not trigger on new regions, example is shown in ablation.
+On Glove, distance computations increase faster due to $REFINE$ regions containing more points and more cracks.
+The median $REFINE$ in Glove100 covers 4.784.78% of the dataset with 512 cracks, while the median $REFINE$ in SIFT1M is only $0.63\%$ of the total points across 64 cracks.
+Nonetheless, the total cost remains amortized over time and is still almost an order of magnitude lower than that of the largest IVF index, while achieving similar final QPS-Recall performance.
+Finally, CRACK is dominated by data movement, while REFINE is mostly compute (Table 2).
+Even though CRACK is infrequent ($0.03\%$ of trace Figure 3), its overhead is not fully captured by distance computations alone.
+This is why we use fixed hardware and cumulative time plots, to compare the total time across all operations, including search, in Figure 4.
+(a) SIFT1M
+(b) SIFT10M
+(c) Last.fm
+(d) Glove50
+(e) Glove100
+(f) Deep10M
+Figure 5. Cumulative distance computations for index build.
+(a) WHERE=OFF
+(b) WHERE=ON
+Figure 6. Effect of removing the “where” control mechanism
+(a) Cumulative Time (OFF)
+(b) Time Breakdown (OFF)
+(c) Cumulative Time (ON)
+(d) Time Breakdown (ON)
+Figure 7. Effect of removing the “when” control mechanism
 5.5. Control Mechanisms Ablation Study
-(a) WHERE=OFF: QPS vs Recall
-(b) WHERE=ON: QPS vs Recall
-Figure 5. Comparison of QPS-Recall, showing effect of removing the ”where” to CRACK and REFINE mechanism
-(a) Cumulative Time (OFF)
-(b) Total Time Breakdown (OFF)
-(c) Cumulative Time (ON)
-(d) Total Time Breakdown (ON)
-Figure 6. Comparison of cumulative time and total time breakdown, showing effect of removing the ”when” to CRACK and REFINE mechanism
-CrackIVF uses two control mechanisms for the CRACK and REFINE build operations, as detailed in Section 4.2 and Section 4.3.
-The first mechanism, controlling where they should be applied, uses a set of heuristic rules.
-The second mechanism, controlling when they are executed, is a cost-budget-based approach combined with a predictive model to balance search and build times.
-This section demonstrates the importance of these controls in achieving optimal performance and additionally highlights the effect of changing the parameter default value $min\_pts=2$, which changes the heuristic rule “Don’t Crack: Rule 1“ (Section 4.2).
-Turning off the control mechanism for “where“: To illustrate the importance and accuracy of the heuristic rules for where to CRACK and REFINE, we switch them OFF and see the effect (Figure 5).
-Specifically, when “WHERE=OFF”, and no heuristic rule is used to decide where a crack or refine should happen, the index defaults to trying to apply these operations whenever there is enough budget for them (the “when“ mechanism is still ON).
-This seemingly random addition of new cracks and local refines leads to a final performance that is no better than the initial state of CrackIVF, before any query has been received (Figure 5(a)).
-On the other hand, when “WHERE=ON“, and our heuristic rules are applied to make decisions, CrackIVF converges to the Pareto optimal QPS-Recall trade-off performance (Figure 5(b)).
-In both examples, we send an equal number of queries to the index and in the same order.
-Turning off the control mechanism for “when“:
-The budgeting mechanism, controlling when build operations are executed, helps balance the time spent on build vs search operations.
-By default, we use the parameter $a=0.5$, i.e., at most 50% of the total time may be spent on $CRACK$ or $REFINE$.
-The effect of turning this mechanism “WHEN=OFF“ is shown in Figure 6(a) and Figure 6(b).
+This section is an ablation study exploring the control mechanisms “where” and “when” to apply $CRACK$ and $REFINE$.
+Turning off the control mechanism for “where”: To illustrate the importance and accuracy of the heuristic rules for where to CRACK and REFINE, we switch them OFF and see the effect (Figure 6).
+Specifically, when “WHERE=OFF”, and no heuristic rule is used to decide where a crack or refine should happen, the index defaults to trying to apply these operations whenever there is enough budget for them, since the “when” mechanism is still ON.
+This seemingly random addition of new cracks leads to a final performance that is only slightly better than the initial state of CrackIVF, before any query has been received (Figure 6(a)).
+On the other hand, when “WHERE=ON”, and our heuristic rules are applied to make decisions, CrackIVF converges to the Pareto optimal QPS-Recall trade-off performance (Figure 6(b)).
+In both examples, we test an equal number of queries and in the same order.
+Turning off the control mechanism for “when”: The budgeting mechanism, controlling when build operations are executed, helps balance the time spent on build vs search operations.
+By default, we use the parameter $\alpha=0.5$, i.e., at most 50% of the total time may be spent on $CRACK$ or $REFINE$.
+The effect of turning this mechanism “WHEN=OFF” is shown in Figure 7(a) and Figure 7(b).
 CrackIVF with the mechanism off defaults to cracking and refining after almost every query, and it only stops when the maximum number of partitions is reached (16,000).
 So even though the cracks and refines happen in regions that require them, they happen disproportionally often compared to the search overhead, leading to a final time spent on build operations that completely dominates the total time.
-On the other hand, when the mechanism is enabled (“WHEN=ON“), CrackIVF is able to efficiently balance the search and build costs (Figure 6(b) and Figure 6(c)).
+On the other hand, when the mechanism is enabled, “WHEN=ON”, CrackIVF is able to efficiently balance the search and build costs (Figure 7(c) and Figure 7(d)).
 These runs are for the same 100k query trace.
-Varying heuristic rule parameters:
-(a) QPS vs Recall ($min\_pts=2$)
-(b) QPS vs Recall ($min\_pts=32$)
-Figure 7. Comparison of QPS vs Recall for different $min\_pts$ values.
-CrackIVF’s heuristic rules use default parameters (Section 4.2).
-To avoid parameter tuning, we did not vary these values across experiments.
-However, testing revealed that the $min\_pts$ parameter can impact performance, dependent on the dataset.
-$min\_pts$ defines the minimum number of points that must be stolen by a new crack for it to be considered a “good“ and subsequently buffered.
-On the GloVe-25 dataset, changing $min\_pts$ from 2 to 32 notably improved the QPS-Recall performance of the final index (Figure 7).
-Originally, we set $min\_pts=2$, a very small value, to ensure we avoid adding empty cracks to the index.
-We assumed that any imbalances could be fixed by future $REFINE$ operations that redistribute points locally, which works for all datasets except for GloVe-25.
-Although we can not make a definite statement, we hypothesize that this can be attributed to the fact that this dataset only has 25-dimensional embeddings.
-This leads to search times that are very low, and subsequently, our build operations budget grows slower.
-Thus, there are too many imbalances introduced with $min\_pts=2$, and a smaller budget for future $REFINE$ operations to correct them.
-By increasing $min\_pts$ to 32, we reduced the total number of local imbalances by avoiding to ever commit excessively small cracks, which in turn reduces the number of $REFINE$ needed to reach a good final performance.
-A second hypothesis comes from closer inspection of the Figures, in which the better performing index converges to 10056 partition vs the default 16000.
-(Figure 7(a), Figure 7(b) ).
-But, we believe because GloVe-25 queries access the entire vector space evenly (Figure 2(b)) and it is a dataset with (1.18M points), we do not think a larger number of partitions hinders performance.
-Nevertheless, we show that parameter values do matter across datasets, so index tuning can become a consideration.
-But for this work, across all the datasets we tested, outside of GloVe-25, the parameter defaults we have chosen have reliable performance.
-Takeaways: Our results show that CrackIVF minimizes the cumulative costs across a large variety of individual datasets tested, but to fully grasp the potential performance benefits in embedding lakes, it’s essential to consider the aggregate across all datasets’ cumulative time savings and as the number of queries varies.
-For systems managing potentially thousands or millions of datasets, all saving are scaled equivalently, and choosing a separate index for each dataset is near-impossible, so the benefit from using CrackIVF is greater, as there is no one-size-fits all solution from the pre-built indexes.
+The parameter controlling the “when” mechanism is $\alpha$.
+The effect of varying $\alpha$ is shown in Figures 8(a) and  8(b).
+If build is 25% of the total time then it is too restrictive.
+On the other hand, if 75% of the total time can go to build operations, a very large number of refines keep triggering.
+Figure 8(b) illustrates the importance of monitoring the new buffered cracks rate and the refine regions in order to converge the index.
+After $\approx$50k queries, no new cracks occur, preventing further index growth.
+Refines also repeatedly trigger in the same regions, as shown by the striated pattern, offering no additional benefit, since the median response time does not improve after the refines.
+This indicates the index has converged.
+We monitor when this state is achieved and then converge the index, avoiding further build operations.
+(a) SIFT1M $\alpha=0.25$
+(b) SIFT1M $\alpha=0.75$
+(c) Deep10M min_pts=32min\_pts=32
+(d) SIFT1M min_pts=32min\_pts=32
+(e) lastfm pts_crack_thr=16pts\_crack\_thr=16
+(f) Glove50 pts_crack_thr=128pts\_crack\_thr=128
+(g) Deep10M cv_max=8cv\_max=8
+(h) SIFT10M cv_max=8cv\_max=8
+(i) SIFT1M size_prctlsize\_prctl OFF
+(j) Glove100 size_prctlsize\_prctl OFF
+Figure 8. Ablation study of parameters.
+Table 4. Parameters defining the Decision Rule Boundaries.
+Parameter Restrict “good candidates” Tested Values
+min_ptsmin\_pts 1, 2, 32
+pts_crack_thrpts\_crack\_thr 16, 64, 128
+cv_maxcv\_max 1, 2, 8
+size_prctlsize\_prctl 1, 10, 25
+Ablation on decision boundary parameters: Each parameter defines a decision boundary (Section 4.2).
+Changing it in one direction, makes the model more restrictive, classifying fewer candidates as good, while in the opposite direction it is more permissive.
+Table 4 indicates the restrictive direction and tested values (default in bold).
+We did multiple runs for each dataset and parameter combination.
+For each parameter we pick the most interesting results, showcased in Figure 8.
+We only vary one parameter at a time, the others are fixed to their default value.
+We found the model to be robust to changing a single parameter.
+Out of the 48 edge-case parameter-dataset combinations, 15 had a noticeable effect on the final index.
+We attribute the robustness to each heuristic involving two decision rules (Section 4.2), thus two parameters.
+When one is changed, the other still provides a reasonable boundary.
+Setting both parameters to restrictive extremes prevents any cracking, while setting both to relaxed extremes is equivalent to WHERE=OFF (Figure 6(a)), where every query is classified as good and the operations occur when the cost budget allows.
+For min_ptsmin\_pts, we have the most consistent effect across all our observations.
+When we set it to a very restrictive value ($32$), the number of good cracks drops significantly, and thus the index converges to a smaller size, leading to lower performance (Figure 8(d)).
+This negatively affected all 6 datasets.
+The pts_crack_thrpts\_crack\_thr, mostly affects the highest skew datasets.
+For example, on Last.fm performance worsens when the parameter is set to a permissive value, as this leads to over-partitioning of the region, as more cracks are allowed in a region (Figure 8(e)).
+But, the above value did not affect any other dataset.
+On the other hand, restricting it negatively affected 2 out of 6 datasets (Figure 8(f)).
+This was the same effect as that of min_ptsmin\_pts.
+When the rule is too restrictive, not enough cracks pass it so index can not fully grow.
+Restricting cv_maxcv\_max can have a positive effect, e.g.
+on SIFT10M (Figure 8(h)) and on DEEP10M (Figure 8(g)).
+This leads us to believe that in some datasets allowing a higher local imbalance, could have beneficial overfitting effects to the query distribution.
+The optimal value for cv_maxcv\_max is not consistent across datasets, two other datasets had worse results at a high value, and our default values provide the most robust results.
+All the values tested for size_prctlsize\_prctl had a negligible effect.
+Thus, to show the benefit of including the decision rule which size_prctlsize\_prctl controls, we switch it off entirely in Figures 8(i) and 8(j).
+Disabling it greatly affects final index.
+While the index grows, and it does not converge to the optimal state, as fewer refines occur.
+Overall, we clearly see dataset-dependent effects which can change the optimal value of each parameter per dataset.
+This leaves room for further tuning and “learned” decision boundaries instead of setting them manually.
+Tuning of parameters exists in any indexing technique, the more you know about your dataset the better you can tune the index for it.
+We did not pursue tuning in this work, as the default values empirically provide very robust performance.
 From our experiments, we make the following suggestion: BruteForce should be used as the default method up until a dataset surpasses 10–100 queries, at which point CrackIVF becomes the optimal choice.
-Even though CrackIVF can start with one partition to simulate a Brute Force approach, it has a slight memory and runtime overhead compared to it, from the setup time and storage of internal object structures and keeping track of the first few partition representatives.
-At scale, for embedding lakes with millions of datasets, even this small overhead can accumulate.
-Therefore, we recommend using CrackIVF once a dataset exceeds 100 queries and is thus likely to receive even more.
+Even though CrackIVF can start with minimal initialization cost, it is more complex and has a slight memory and runtime overhead compared to BruteForce due to the setup time and storage of internal object structures and keeping track of the first few partition representatives.
+At scale, in embedding lakes even a small overhead can accumulate.
+We recommend using CrackIVF once a dataset exceeds 100 queries and is likely to receive more.
 6. Related Work
-Vector Databases: Vector Data Management Systems (VDBMS) (Wang et al., 2021; Weaviate, 2025; Pinecone, 2025; Zhang et al., 2023; Chen et al., 2024; Guo et al., 2022), are systems that support efficient ANN search, and have gained significant popularity in both research and industry.
-This coincides with the increase in popularity of Large Language Models (LLMs) (Brown et al., 2020), that use Retrieval-Augmented Generation (RAG) techniques (Lewis et al., 2020; Asai et al., 2023), to extend their capabilities beyond what they have been trained on, in order to answer queries based on external data sources.
-Vector Database Management System applications extend beyond the above use-case, as they are also used as platforms for semantic search (Salton et al., 1975; Mitra et al., 2018), recommendation systems (Li et al., 2017) and product search in e-commerce (Van Gysel et al., 2016).
-Given the importance of these operations, many existing data management systems are incorporating vector search capabilities (Corporation, 2025; pgvector contributors, 2025).
-Many of these systems also support extensions to dense vector ANN search, through specialized methods, allowing for specific query types such as filtered search (Gollapudi et al., 2023), sparse vector search (Formal et al., 2021), and hybrid search (Vespa.ai, 2024), which combines vector search with more traditional information retrieval techniques like BM-25 (Robertson et al., 2009).
-A recent survey of vector database management systems is found in (Pan et al., 2024).
-Vector Search Indexes: Approximate Nearest Neighbor (ANN) search is a well-studied field with decades of research, resulting in a large number of indexes.
-Broadly, the two dominant categories of ANN vector search indexes are partition-based and graph-based indices.
-Inverted index (IVF) based methods, which fall under partition based category, include SCANN and SOAR (Guo et al., 2020; Sun et al., 2024), which identify key mathematical properties of ANNS and use them to improve IVF index efficiency, SPANN (Chen et al., 2021), which shows how to leverage disk-based storage for search over billion-scale datasets along with supporting updates (Xu et al., 2023), and FAISS-IVF (Douze et al., 2024), an in-memory implementation utilizing GPUs (Johnson et al., 2019) and quantization techniques like PQ (Sivic and Zisserman, 2003; Jegou et al., 2010; Ge et al., 2013).
-In this work, we focus on IVF methods, and our cracking based approach could be extended for the above implementations.
-IVF indexes offer the best trade-off between index build time, search performance, and memory overhead, all of which are critical to balance for potential large scale deployments in embedding data lakes.
-Hash based methods (Andoni et al., 2015; Indyk and Motwani, 1998; Datar et al., 2004; Jain et al., 2008; Wang et al., 2017) and tree based methods (Muja and Lowe, 2014; Liu et al., 2004; Wang et al., 2013) also broadly fall under the partition-based index category.
+Vector Databases: Vector Data Management Systems (VDBMS) (Wang et al., 2021; Weaviate, 2025; Pinecone, 2025; Zhang et al., 2023; Chen et al., 2024; Guo et al., 2022; Deng et al., 2025; Kuffo et al., 2025; Xu et al., 2025; Cai et al., 2024) support efficient ANN search, and have gained significant popularity.
+This coincides with the increase in use of Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Asai et al., 2023), to extend the capabilities of LLMs beyond what they have been trained on.
+VDBMS applications also include semantic search (Salton et al., 1975; Mitra et al., 2018), recommendation systems (Li et al., 2017) and product search (Van Gysel et al., 2016).
+Many relational engines are incorporating vector search (Corporation, 2025; pgvector contributors, 2025).
+Many of these systems support extensions to dense vector ANN search, allowing for specific query types such as filtered search (Gollapudi et al., 2023), sparse vector search (Formal et al., 2021), and hybrid search (Vespa.ai, 2024).
+A recent survey covering vector database management systems can be found in (Pan et al., 2024).
+Vector Search Indexes: Approximate Nearest Neighbor (ANN) search has been well-studied.
+The two dominant categories of ANN indexes are partition-based and graph-based indices.
+Inverted file index (IVF) is a partition-based index and includes many different implementations.
+ScaNN (Guo et al., 2020) and SOAR (Sun et al., 2024) identify key mathematical properties of ANNS and use them to improve index efficiency.
+SPANN (Chen et al., 2021) uses disk-based storage to search over billion-scale datasets and supports updates (Xu et al., 2023).
+FAISS-IVF (Douze et al., 2024) provides an in-memory implementation IVF which can utilize GPUs (Johnson et al., 2019).
+In this work we show how IVF indexes can be extended with cracking, and our approach is directly applicable to the above.
+We focus on IVF indexes as they offer the best trade-off between index build time, search performance, and memory overhead, which is critical to minimize overhead in large scale deployments on embedding data lakes, as covered in Section 3.
+The cracking paradigm can be generalized to other ANN indexes as well.
+With AV-tree(Lampropoulos et al., 2023) the authors show how to extend a tree-based index with cracking to supports k-NN queries targeting ENN search.
+Hash-based methods (Andoni et al., 2015; Indyk and Motwani, 1998; Datar et al., 2004; Jain et al., 2008; Wang et al., 2017) and tree-based methods (Muja and Lowe, 2014; Liu et al., 2004; Wang et al., 2013) also broadly fall in the partition-based category.
+Each index has a unique construction method and data structures.
+Adopting cracking ideas for them would require solving unique challenges.
 Finally, graph-based methods (Malkov and Yashunin, 2018; Malkov et al., 2014; Wang et al., 2012; Dong et al., 2011; Wang et al., 2013; Hajebi et al., 2011) currently deliver the state-of-the-art performance in terms of the Queries Per Second vs.
 Recall trade-off, but this comes at high memory and index construction cost.
-Thus, graph based indexes are for now ill-suited for scenarios where minimizing the index construction cost and low memory overhead is a important, for example in embedding lakes where a very large number of indexes is expected to be constructed.
-Adaptive indexes for specialized data types: Adaptive indexing techniques, such as Dumpy (Wang et al., 2023, 2024) and ADS (Zoumpatianos et al., 2014), have demonstrated significant effectiveness in data series similarity search.
-These indexes are specifically designed for time-series data and include methods like Symbolic Aggregate Approximation (SAX) (Lin et al., 2003) and indexable SAX (iSAX) (Shieh and Keogh, 2008).
-In our work, we focus on ANN indexing methods that are the standard used by RAG systems, and due to their broad applicability across various data modalities which lack sequential relationships.
-That said, specialized, adaptive and incrementally constructed indexes are relevant candidates for embedding lake deployments, particularly for the subset of the datasets with inherent sequential dependencies.
-This includes energy, weather and financial data, as well as audio and video recordings (Raza et al., 2015; Mirylenka et al., 2016; Shasha, 1999; Kashino et al., 1999).
+Thus, graph-based indexes are ill-suited when the goal is minimizing the index construction cost and memory overhead.
+Cracking is applicable here and could push search performance even further.
+For example, future research can show how to dynamically change the connections in the graphs based on the arriving query distribution.
+Adaptive indexes for specialized data types: Adaptive indexing techniques, such as Dumpy (Wang et al., 2023, 2024) and ADS (Zoumpatianos et al., 2014) are effective in data series similarity search.
+They are specifically designed for time-series data and include methods like Symbolic Aggregate Approximation (SAX) (Lin et al., 2003) and indexable SAX (iSAX) (Shieh and Keogh, 2008).
+We focus on ANN indexing methods that are the standard used by RAG systems, and due to their broad applicability across various data modalities which lack sequential relationships.
+Specialized, adaptive and incrementally constructed indexes are relevant candidates for embedding lake deployments, particularly for the subset of the datasets with inherent sequential dependencies such as energy, weather, financial data, audio, and video (Raza et al., 2015; Mirylenka et al., 2016; Shasha, 1999; Kashino et al., 1999).
 ANN index maintenance: Work on index maintenance focuses on mitigating performance degradation caused by data distribution shifts due to data updates (insertions and deletions).
-DeDrift (Baranchuk et al., 2023) handles data or query distribution shifts and identifies small and large partitions to re-cluster together, but it keeps the total number of partitions fixed.
-LIRE (Xu et al., 2023) performs localized split and merge operations when partitions grow too large or become too small after insertions and deletions.
-Ada-IVF (Mohoney et al., 2023) monitors partition size changes from updates and the partition access distribution from the queries.
-It combines this information to prioritize the maintenance operations that will improve index efficiency.
-Similar to us, these approaches demonstrate the effectiveness of localized maintenance operations for IVF indexes and the advantages of adapting to query distribution.
-However, they cannot be applied directly for our usage, which is incremental index construction as the number of queries increases.
-These works all trigger maintenance operations directly or implicitly from incoming updates (insertions/deletions).
-Thus they will never trigger in our static setting, whereas we do since we follow the database cracking paradigm, that states “index maintenance should be a byproduct of query processing, not of updates” (Idreos et al., 2007), and CrackIVF operations are triggered solely from the queries.
-7. Conclusion and Future Work
+DeDrift (Baranchuk et al., 2023) handles data or query distribution shifts keeping the total number of partitions fixed.
+LIRE (Xu et al., 2023) performs localized split and merge operations when partitions change after insertions/deletions.
+Ada-IVF (Mohoney et al., 2024) monitors partition size changes from updates and the partition access distribution from the queries.
+It combines this information to prioritize the maintenance operations.
+Finally, Quake (Mohoney et al., 2025) is an adaptive index for dynamic workloads with changing access patterns and heavy updates.
+They use a cost model to decide on maintenance operations, and dynamically set the nprobe parameter.
+Similar to us, these approaches demonstrate the effectiveness of localized operations for IVF indexes and the advantages of adapting to query distribution.
+However, they cannot be applied directly for our setting of incremental index construction to minimize cumulative time as the number of queries increases.
+These works trigger maintenance operations from incoming updates, or focus on distribution shits.
+Our focus is on the static setting, and we follow the database cracking paradigm: “index maintenance should be a byproduct of query processing, not of updates” (Idreos et al., 2007).
+Thus, CrackIVF operations are triggered solely from the queries.
+7. Conclusion
 We introduced CrackIVF, a cracking-based IVF index for Approximate Nearest Neighbor Search (ANNS), aimed to be used with RAG systems over embedding data lakes.
@@
 This makes it particularly effective for cold data, infrequently accessed datasets, and rapid access to unseen data.
-As future work, we want to adapt CrackIVF to handle query and data distribution shifts and identify reliable methods to dynamically set parameters across diverse datasets.
-8. Acknowledgments
+Acknowledgements.
 We would like to thank AMD Research for the generous donation of the equipment in the ETHZ-AMD Heterogeneous Accelerated Compute Cluster (HACC) (https://systems.ethz.ch/research/data-processing-on-modern-hardware/hacc.html) that was used to run the experiments reported in this paper.